### v1

# Sparse Subnetwork Enhancement for Underrepresented Languages in Large Language Models

Abstract
Large language models exhibit uneven performance across languages, with substantial gaps between high- and low-resource languages. We present a framework for enhancing monolingual capabilities of LLMs in underrepresented languages while preserving their general-purpose performance through targeted fine-tuning of language-specific subnetworks. Our approach identifies language-specific neurons using Language Activation Probability Entropy and fine-tunes only the weights associated with these neurons, a dedicated subnetwork, on target-language data. Experiments on Llama-3.1-8B and Mistral-Nemo-12B across 12 mid- and low-resource languages demonstrate that our method consistently outperforms full fine-tuning, FFN-only fine-tuning, LoRA adaptation, and random subset fine-tuning baselines while efficiently updating only up to 1% of model parameters. Beyond performance improvements, we observe enhanced favorable training dynamics, cross-lingual representational alignment, and systematic weight update changes. To facilitate future research, we release language-specific neuron identifications for over 100 languages as well as our adaptation pipeline, offering a cost-effective pathway for adapting state-of-the-art models to underrepresented languages.
Sparse Subnetwork Enhancement for Underrepresented Languages
in Large Language Models
Daniil Gurgurov1,2  Josef van Genabith1,2  Simon Ostermann1,2,3 1Saarland University 2German Research Center for Artificial Intelligence (DFKI) 3Centre for European Research in Trusted AI (CERTAIN) daniil.gurgurov@dfki.de
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks and languages, yet their performance remains unevenly distributed, with substantial gaps between high-resource and low-resource languages Joshi et al. (2020); Huang et al. (2023). While scaling training data and model parameters has improved multilingual performance, this approach is still often impractical and ineffective for resource-constrained scenarios involving hundreds of languages Robinson et al. (2023).
Figure 1: Sparse subnetwork adaptation method: attention blocks remain frozen (grey) while selected language-specific FFN neurons and their associated weights are fine-tuned (red). This approach enables parameter-efficient adaptation for underrepresented languages while preserving the model’s general knowledge.
Recent advances in mechanistic interpretability have revealed that language models develop specialized internal structures for different languages, with distinct neurons and attention patterns activated for specific linguistic contexts Tang et al. (2024); Zhao et al. (2024); Kojima et al. (2024). This observation suggests a promising alternative to full model retraining: identifying and selectively enhancing language-specific components while preserving general-purpose capabilities.
Current approaches to multilingual adaptation face a fundamental trade-off between target-language performance and retention of general capabilities. Full fine-tuning often leads to catastrophic forgetting French (1999), while parameter-efficient methods like adapters Houlsby et al. (2019a), LoRA Hu et al. (2021), and soft prompts Vykopal et al. (2025) may not fully exploit the model’s existing multilingual structure. Existing neuron-level approaches have shown promise but have primarily focused on task-specific adaptation rather than systematic language enhancement, and they have often fallen short in preserving existing knowledge Zhu et al. (2024); Xu et al. (2024); Mondal et al. (2025).
In this work, we present a framework for enhancing monolingual capabilities of LLMs through targeted fine-tuning of language-specific subnetworks within feed-forward network (FFN) components. Our approach leverages Language Activation Probability Entropy (LAPE) Tang et al. (2024) to identify neurons that exhibit strong language-specific activation patterns within Llama-3.1-8B and Mistral-Nemo-12B, then fine-tunes only the weights associated with these neurons using target-language data. This sparse fine-tuning updates up to 1% of model parameters while achieving substantial improvements in target-language performance and preserving general capabilities.
We make the following contributions:
  • We introduce a systematic framework for efficient and effective language subnetwork identification and enhancement that preserves general-purpose performance while significantly improving target-language capabilities.
  • We demonstrate the effectiveness of our approach across 12 mid- and low-resource languages, showing consistent improvements over full fine-tuning, FFN-only fine-tuning, LoRA fine-tuning, and random subset fine-tuning baselines.
  • We show that target-language performance can be improved by increasing the amount of training data where possible, laying the foundation for optimal adaptation strategies across different languages.
  • Through activation analysis, we show that targeted subnetwork fine-tuning improves cross-lingual alignment, while weight-change analysis reveals that its effects are concentrated in the later down-projection matrices of FFN layers, highlighting their central role in adapting to new languages.
  • We release language-specific neuron identifications for over 100 languages, providing a foundation for future research in low-resource language adaptation. [^1] https://github.com/d-gurgurov/Language-Subnetwork-Enhancement-LLMs.
This work represents a step toward democratizing high-quality language model capabilities across diverse languages. By providing both a practical framework and comprehensive resources, we aim to lower the barrier for adapting state-of-the-art models to underrepresented languages.
2 Related Work
Prior work on adapting multilingual language models to underrepresented languages can be grouped into three main directions: full fine-tuning, adapter-based approaches, and more recent neuron- or head-level methods.
2.1 Full Fine-Tuning Approaches
Full fine-tuning is the most direct strategy for language adaptation. Post-training on related languages has been used to mitigate overfitting Neubig and Hu (2018), while domain-adaptive fine-tuning improved contextualized models on specialized corpora Han and Eisenstein (2019). Language-specific fine-tuning on monolingual corpora Gururangan et al. (2020); Chau et al. (2020), transliteration-based adaptation Muller et al. (2021), and even small corpora such as Bible translations Ebrahimi and Kann (2021) have boosted performance on tagging, parsing, and NER. However, full fine-tuning is costly, data-hungry, and prone to catastrophic forgetting French (1999), motivating parameter-efficient alternatives.
2.2 Adapter-Based Approaches
Adapters Houlsby et al. (2019b) introduce small trainable modules into frozen backbones, reducing computational cost Strubell et al. (2019) and data requirements Faisal and Anastasopoulos (2022) while avoiding forgetting French (1999). MAD-X Pfeiffer et al. (2020), UDapter Üstün et al. (2020), and MAD-G Ansell et al. (2021) extend adapters with task/language modularity and typological signals, while other work explores hierarchical designs Faisal and Anastasopoulos (2022), language balancing Lee et al. (2022); Parović et al. (2022), and inference-aware training Parović et al. (2023). Recent studies show adapter tuning can outperform continued pre-training for adaptation Yong et al. (2023); Gurgurov et al. (2025b).
2.3 Neuron-Level Approaches
Recent work has explored fine-grained adaptation by targeting individual neurons or attention heads. Zhu et al. (2024) identify language-aware neurons in FFNs using Taylor scores Xie et al. (2021) and fine-tune only the relevant subsets using parallel data for machine translation, outperforming both adapters and full fine-tuning in high-resource settings. Similarly, Zhao et al. (2024) adapt $\sim$0.1% of neurons in FFNs and attention heads identified via neuron ablations, observing initial gains for a subset of languages but rapid saturation with additional data. Xu et al. (2024) propose NeFT, which compares pre- and post-fine-tuning representations to locate task-sensitive neurons, and show that fine-tuning just 5–10% of them consistently surpasses LoRA Hu et al. (2021) on translation and summarization. Mondal et al. (2025) apply LAPE Tang et al. (2024) for neuron identification and manipulation, but report limited improvements in cross-lingual transfer. Their setup, however, differs from ours in that they restrict analysis to only three relatively high-resource languages and adapt LoRA modules on top of FFNs rather than directly fine-tuning neurons inside the model, which makes their findings not directly comparable to ours. Complementary work shows that only a few attention heads encode translation mechanisms, and selectively adapting these heads matches full fine-tuning while preserving general abilities Zhang et al. (2025).
2.4 Towards Targeted Adaptation
Neuron- and head-level methods demonstrate that precise adaptation can rival or exceed large-scale fine-tuning, yet prior work has emphasized task-specific gains rather than systematic language adaptation. In particular, it remains unclear to what extent language-specific information can be injected into carefully chosen subnetworks for underrepresented languages without harming general performance. Our work addresses this gap by leveraging LAPE, an information-theoretic method that has recently shown strong potential for precise language neuron identification Tang et al. (2024) and manipulation Gurgurov et al. (2025a), to identify language-specific subnetworks and fine-tune their associated weights on target-language data.
3 Methodology
Our approach consists of two components: (i) identifying language-sensitive neurons, and (ii) selectively fine-tuning the corresponding subnetwork (see Figure 1). Together, these steps allow us to adapt LLMs to underrepresented languages while minimizing changes to unrelated parameters.
3.1 Neuron Identification
We leverage Language Activation Probability Entropy (LAPE) Tang et al. (2024) to locate neurons that exhibit language-specific behavior. We focus on neurons within FFN components, since prior work has shown that they frequently encode rich factual knowledge Geva et al. (2021); Dai et al. (2021); Wang et al. (2022) and multilingual information Zhao et al. (2024).
For each neuron $j$ in layer $i$, we compute the probability of activation under inputs from language $k$:
$p^{k}_{i,j}=\mathbb{E}_{x\sim\mathcal{D}_{k}}\Big[\mathbf{1}\!\left(\phi(h_{i}(x)W^{i}_{1})_{j}>0\right)\Big],$
where $h_{i}(x)$ is the hidden state entering layer $i$, $W^{i}_{1}$ is the input projection of the feed-forward network, $\phi$ is the activation function, and $\mathbf{1}$ is the indicator function.
The vector of probabilities across all languages $k=1,\dots,L$ is normalized:
$\tilde{p}_{i,j}^{k}=\frac{p^{k}_{i,j}}{\sum_{\ell=1}^{L}p^{\ell}_{i,j}}.$
The LAPE score is then given by Shannon entropy Shannon and Weaver (1998):
$\text{LAPE}_{i,j}=-\sum_{k=1}^{L}\tilde{p}_{i,j}^{k}\log\tilde{p}_{i,j}^{k}.$
We select neurons with the lowest $K\%$ of LAPE scores, indicating truly language-specific behavior due to their low-entropy activations, subject to two conditions:
1. $\max_{k}p^{k}_{i,j}\geq\tau_{\text{activity}}$: a neuron must reach at least $\tau_{\text{activity}}$ in some language; otherwise, uniformly weak neurons are discarded.
2. $\exists k:p^{k}_{i,j}\geq\tau_{\text{selectivity}}$: a neuron is kept only if its activation for at least one language exceeds a strong threshold $\tau_{\text{selectivity}}$, ensuring clear language preference.
The resulting subset forms a language-sensitive subnetwork $\mathcal{S}^{k}$.
3.2 Subnetwork Fine-Tuning
Given the identified subnetwork $\mathcal{S}^{k}$, we restrict parameter updates to the weights associated with these neurons. Let $\theta$ denote the full set of model parameters, and $\theta_{\mathcal{S}^{k}}\subset\theta$ the parameters belonging to the selected neurons.
Fine-tuning is then performed by minimizing the loss $\mathcal{L}$ on target-language data $\mathcal{D}_{\text{target}}$:
$\min_{\theta_{\mathcal{S}^{k}}}\;\mathbb{E}_{(x,y)\sim\mathcal{D}_{\text{target}}}\;\mathcal{L}\big(f(x;\theta),y\big),$
while keeping all parameters $\theta\setminus\theta_{\mathcal{S}^{k}}$ fixed.
This optimization aims at adapting only the language-relevant subnetwork while preserving the general-purpose behavior of the model.
4 Experimental Setup
Below, we describe the details of our training setup, the evaluation tasks chosen to assess model performance, and the baselines against which we compare.
4.1 Training Setup
We conduct experiments with Llama-3.1-8B-base and Mistral-Nemo-12B-base Grattafiori et al. (2024), using GlotCC Kargaran et al. (2024) as the source of target-language training data. By default, each language subnetwork is trained on up to 100M tokens in the target language, with additional experiments varying the training size between 50M and 200M tokens for a subset of languages. We focus on 12 underrepresented languages Joshi et al. (2020), reported in Table 1.
Subnetworks are identified for each language using 100MB of GlotCC data, with thresholds $K=5\%$ and activity thresholds of 0.95, following Gurgurov et al. (2025a). Each subnetwork is fine-tuned for one epoch with batch size 2, a learning rate of $1\times 10^{-4}$, and weight decay of 0.01. We use AdamW Loshchilov and Hutter (2019) with gradient clipping Mikolov et al. (2012) for stability. Validation is performed every 500 steps, and the best checkpoint is selected by validation loss on the corresponding Flores-200 subset Costa-Jussà et al. (2022).
Lang. # Neurons # Trainable Param-s %
Afrikaans 2581 31.7M 0.39
Welsh 6020 74.0M 0.92
Estonian 3204 39.4M 0.49
Icelandic 5033 61.8M 0.77
Georgian 1150 14.1M 0.18
Lithuanian 2978 36.6M 0.46
Latvian 3212 39.5M 0.49
Macedonian 1743 21.4M 0.27
Maltese 3015 37.0M 0.46
Nepali 4625 56.8M 0.71
Slovak 1940 23.8M 0.30
Slovenian 2046 25.1M 0.31
Table 1: Neuron, parameter counts for different language subnetworks, and percentages of model being trained within Llama-3.1-8B. The statistics for Mistral-Nemo-12B are available in Appendix A.
4.2 Evaluation Tasks
We evaluate our method along two dimensions:
Target-language performance. We assess translation quality using the "valtest" splits of Flores-200 Costa-Jussà et al. (2022) and the "test" splits of OPUS-100 Tiedemann (2012), reporting results in Sentence BLEU Post (2018). For comprehension and classification, we evaluate on Belebele Bandarkar et al. (2024) and SIB-200 Adelani et al. (2023), using accuracy as the primary metric.
General capabilities. To measure the impact of adaptation on overall model performance, we report accuracy on several standard benchmarks before and after fine-tuning: MMLU Hendrycks et al. (2021), Hellaswag Zellers et al. (2019), PIQA Bisk et al. (2019), Winogrande Sakaguchi et al. (2019), and ARC (Easy and Challenge) Clark et al. (2018).
For all tasks, we use deterministic generation with zero temperature to eliminate variance introduced by stochastic decoding. The prompts and hyperparameters specific to each evaluation are listed in Appendix C.2.
Figure 2: Neuron distributions over the layers of Llama-3.1-8B for all 12 languages. The distributions for Mistral-Nemo-12B are in Appendix A.
Metric Base Model Full FT FFN-Only FT LoRA Random Neurons Target Neurons
Target Language Performance
FLORES tg→en 28.82 11.44 25.06 28.00 30.95 31.37
FLORES en→tg 11.62 9.29 13.70 15.60 15.84 16.57
BELEBELE 66.52 21.22 38.35 56.61 59.82 66.84
SIB200 40.89 11.23 21.65 33.37 34.07 44.94
OPUS tg→en 19.59 10.67 20.01 22.68 24.52 24.34
OPUS en→tg 13.32 7.71 13.04 16.33 16.93 17.23
Avg. 29.56 11.88 21.81 28.41 29.89 33.02
General Capabilities
ARC challenge 70.90 20.40 32.94 59.45 63.69 70.62
ARC easy 85.79 23.86 46.77 77.12 80.83 84.72
Hellaswag 40.38 13.49 22.56 35.02 34.90 40.19
PIQA 71.87 34.84 48.19 64.68 69.69 72.55
Winogrande 51.40 38.15 48.34 50.76 51.24 52.41
MMLU 58.06 22.34 32.59 49.59 52.45 57.52
Avg. 63.07 25.51 38.57 56.10 58.80 63.00
Table 2: Average results for enhancing language subnetworks within Llama-3.1-8B, evaluated on both monolingual capabilities and general performance. All models were trained on 100M tokens in the target language. The downstream results for the individual languages are in Appendix C.
4.3 Baselines
We compare our sparse fine-tuning approach against four baselines:
  • (i) full model fine-tuning,
  • (ii) fine-tuning only the FFN components,
  • (iii) fine-tuning a random subset of FFN neurons of the same size as the LAPE-selected subnetwork, and
  • (iv) fine-tuning a LoRA Hu et al. (2021) applied to gate_proj and gate_up, configured to match the average parameter count used for most languages, with rank set to 62 and alpha set to 32, resulting in 57M trainable parameters (0.7%).
For full-model and FFN-only fine-tuning, we set the learning rate to $1\times 10^{-5}$ to account for the larger number of trainable parameters, while keeping all other hyperparameters identical to those of our method. For LoRA and the random-subset baseline, we fine-tune with exactly the same hyperparameters as our method.
5 Experimental Results
In this section, we present our experimental findings. We first report the results of language-specific subnetwork identification, analyzing the extent to which distinct subnetworks emerge across languages. We then evaluate the effectiveness of subnetwork fine-tuning, highlighting its impact on performance relative to baselines.
5.1 Subnetwork Identification
Table 1 summarizes the number of language-specific neurons, the corresponding trainable parameters, and the fraction of the model trained for each language-specific subnetwork within Llama-3.1-8B. The table shows substantial variation across languages: for instance, Welsh and Nepali subnetworks are relatively large, covering over 0.7% of the model parameters, whereas Georgian and Macedonian subnetworks are much smaller, accounting for less than 0.3%. This variation might reflect differences in language complexity and the amount of language-specific signal captured by GlotCC. The subnetwork sizes for Mistral-Nemo-12B are in Appendix A.
Figures 2 and 11 depict the distribution of the subnetwork neurons across layers for all 12 languages. We observe that most language-specific neurons are concentrated in the upper layers, and related languages exhibit moderate overlap in their neuron sets (such as Slovak and Slovenian, or Latvian and Lithuanian–see Appendix A for details). These patterns align with prior findings from Gurgurov et al. (2025a). Overall, the weights associated with these subnetworks account for roughly 0.2% to 1% of the total model parameters, highlighting the efficiency of this approach.
5.2 Subnetwork Fine-tuning
5.2.1 Average Task Performance
Table 2 reports average downstream task performance across the 12 target languages for different fine-tuning configurations, along with results on general-purpose benchmarks for Llama-3.1-8B. Targeted sparse fine-tuning of language-specific subnetworks consistently yields the largest gains on target-language tasks, outperforming full-model, FFN-only, LoRA, and random-subset baselines, while maintaining general capabilities as measured by MMLU and commonsense tasks. For example, on FLORES, targeted subnetworks achieve average improvements of up to 5 points, while preserving or slightly improving performance on general benchmarks such as MMLU and Winogrande. This contrasts with the findings of Mondal et al. (2025), who also employed LAPE for neuron identification but restricted their analysis to only three relatively high-resource languages and adapted LoRA modules on top of FFNs rather than directly fine-tuning neurons inside the model, leading to more limited improvements in cross-lingual transfer. By contrast, full-model and FFN-only fine-tuning not only fail to improve target-language performance but also substantially degrade general capabilities. The random neuron-subset and LoRA baselines show intermediate behavior, with modest target-language gains but noticeable reductions in overall performance, underscoring the advantage of precisely selecting language-specific subnetworks for sparse adaptation. Results for individual languages for both Llama-3.1-8B and Mistral-Nemo-12B are provided in Appendix C.
(a) Llama-3.1-8B
(b) Mistral-Nemo-12B
Figure 3: Average improvement scores across all tasks for the target fine-tuning for each of the 12 languages.
5.2.2 Language-Wise Variation
As shown in Figure 3, the magnitude of improvement varies considerably across languages. Languages such as Afrikaans, Maltese, and Welsh benefit the most from targeted fine-tuning, with average gains exceeding 5 points across all tasks, whereas languages such as Latvian and Macedonian see only marginal improvements. A likely explanation is that higher-resource languages are already well represented in pre-training, limiting the marginal utility of additional fine-tuning data. In some cases, the model may even have encountered overlapping material during pre-training, further reducing the benefit of adaptation Lee et al. (2021). By contrast, lower-resource languages, on which the base model underperforms, gain substantially more from precise subnetwork adaptation. The only decreases we observe are for Georgian with Llama and Slovenian and Nepali with Mistral. Since both models show improvements on these languages overall (i.e., Mistral for Georgian and Llama for Slovenian and Nepali), and improvements are also evident on individual tasks (here we report averages), the gains for these languages are clearly possible but may not materialize due to the relatively small number of subnetwork parameters allocated to them. Finally, we note that some languages–particularly highly morphological ones or those with different scripts–do improve under our methods, but less than languages such as Afrikaans and Maltese. This may indicate that they require larger or smaller subnetworks (i.e., higher or lower K%). We leave the systematic exploration of optimal subnetwork sizes for different languages to future work.
Figure 4: Validation loss curves for different fine-tuning configurations of Llama-3.1-8B on Afrikaans (all trained on 100M tokens). The plots for the other tested languages are in Appendix B.
Figure 5: Overall improvement scores against various data sizes for a subset of 5 languages when sparsely fine-tuning Llama-3.1-8B.
(a) Down Projection
(b) Up Projection
(c) Gate Projection
Figure 6: Per-layer changes in weight values for three considered weight types for Nepali in Llama-3.1-8B. The plots for the other 11 languages follow the same patterns and will be made available in our GitHub repository.
(a) Original Llama-3.1-8B
(b) Fine-tuned (100M tokens)
(c) Average cosine similarity
Figure 7: Cosine similarity for post-FFN hidden state activations of the original and sparsely fine-tuned Llama-3.1-8B between Maltese (left, middle) and average across target languages (right). Full results in Appendix D.
5.2.3 Training Dynamics
Figure 4 illustrates the training dynamics for Afrikaans. Sparse fine-tuning of target-language subnetworks generally converges faster and reaches lower validation loss than random-subset, FFN-only, full-model, or LoRA fine-tuning. While full-model and FFN-only fine-tuning occasionally achieve lower loss on the target language (see Appendix B) due to the greater capacity of the updated networks, this is often accompanied by pronounced overfitting and a substantial drop in general capabilities.
5.2.4 Effect of Training Data Size
Figure 5 further examines the effect of varying training data size (50M, 100M, 150M, and 200M tokens) for a subset of seven languages (Macedonian, Latvian, Nepali, Slovak, Afrikaans, Maltese, and Lithuanian) using Llama-3.1-8B. We observe relatively consistent improvement trends with increasing data across most tested languages, irrespective of subnetwork size, suggesting that additional data could further enhance performance. This differs from the findings of Zhao et al. (2024), who fine-tune an ablation-identified subset of neurons using between 100 and 800 documents–likely comprising fewer tokens–and report saturation after 400 documents. However, for Lithuanian we similarly observe a consistent decrease with larger training token sizes, which may indicate that the additional data is not helpful Lee et al. (2021) or that the capacity of the identified subnetwork is insufficient to effectively utilize it. Due to computational constraints and the limited availability of data for some languages (e.g., Maltese and Afrikaans, with at most 100–150M tokens), we leave the exploration of larger-scale training to future work.
6 Effects of Sparse Fine-Tuning
In this section, we present additional analyses focusing on two aspects: the changes in model weights induced by our sparse fine-tuning approach and the evolution of representation similarities before and after fine-tuning.
6.1 Weight Change Analysis
Our analysis of the weight changes during fine-tuning reveals a notable asymmetry between the up/gate projections and the down projection. As shown in Figure 6, the distribution of weight changes for the gate_proj and up_proj remains relatively consistent across all layers. The small size of the box plots indicates that most weight changes are exceptionally small and tightly clustered around zero. This suggests that the fine-tuning process is highly precise, making only minor adjustments to most weights. In contrast, the weight changes in the down_proj show an increase in magnitude across all layers, with especially large shifts in the later layers. This suggests that individual weights undergo more substantial alterations there, aligning with the greater prominence of language-specific neurons in those parts of the model. This finding suggests a functional distinction in how different weight components of the FFN layers contribute to language fine-tuning.
This asymmetry can be explained by the distinct roles of these projections within the SwiGLU-based architecture Shazeer (2020). The up_proj and gate_proj work together to create a gated feature space, refining the model’s ability to select and activate an appropriate subset of existing features for the new language. Updating these weights primarily involves slightly re-weighting established feature-extraction capabilities rather than fundamentally altering the feature space itself. Conversely, the down_proj is responsible for an important integrative step, taking the activated features and synthesizing a new representation. Our observation that its weight changes are substantially larger in later layers indicates this is where the most significant language-specific re-wiring occurs, as these deeper layers learn to reorganize high-level concepts to adapt to the new language.
6.2 Representation Analysis
Targeted neuron fine-tuning not only improves performance in the target language but also enhances cross-lingual alignment. To investigate this, we examine whether parallel sentences become closer in embedding space after adaptation. Figure 7 presents cosine similarities between post-layer hidden state activations of Flores-200 sentences in Maltese and several high-resource languages (Arabic, Italian, English, French, German, Spanish, Russian, Chinese, Japanese) across layers. We observe a consistent increase in similarity for all languages after adaptation, especially in the later layers, as the early layers already exhibit relatively high similarity.
The average results across all languages in Figure 7(c) show that targeted sparse fine-tuning achieves the strongest cross-lingual alignment, with an average cosine similarity of 90.47 between parallel sentences. This represents a substantial improvement over all considered fine-tuning baselines. These findings indicate that selectively updating neurons identified as important for the target language produces more aligned representations across languages, suggesting that sparse adaptation can improve cross-lingual generalization without compromising overall model structure. Consistent with these results, Sundar et al. (2025) also report improved cross-lingual alignment when manipulating language-relevant activations.
7 Conclusion
We presented a framework for enhancing the capabilities of LLMs in underrepresented languages through targeted fine-tuning of language-specific subnetworks. By identifying neurons with strong language selectivity using LAPE and adapting only their associated weights, our approach achieves substantial gains in low- and mid-resource languages while preserving general-purpose performance and updating less than 1% of parameters. Beyond improved accuracy, we showed that targeted adaptation enhances cross-lingual alignment and yields favorable training dynamics. Our release of language-specific neuron identifications for over 100 languages provides a scalable and cost-effective resource for extending state-of-the-art models to underrepresented languages, helping narrow the gap in global language coverage.
Limitations
While our framework improves performance across diverse languages, it relies on accurate identification of language-specific neurons, which can be sensitive to dataset quality and language coverage. Our experiments primarily focus on mid- and low-resource languages, leaving open the question of how the method scales to extremely low-data settings or to languages that are highly typologically distant. Furthermore, we do not explore varying the capacity of identified language-specific subnetworks through the parameter K%. Understanding how many parameters are needed to effectively capture language-specific knowledge, and how much data is required to leverage them, remains an important direction for future work. Such analyses, conducted at the per-language or per-language-family level, have significant potential.
Acknowledgments
This research was supported by the German Federal Ministry of Research, Technology and Space (BMFTR) as part of the project TRAILS (01IW24005) and lorAI - Low Resource Artificial Intelligence, a project funded by the European Union under GA No.101136646.
Appendix
Appendix A Language Neuron Overlap and Individual Distributions
Figure 8: Language neuron overlap for all 12 languages for Llama-3.1-8B.
Figure 9: Language neuron overlap for all 12 languages for Mistral-Nemo-12B.
Figure 10: Neuron distributions over the layers of Mistral-Nemo-12B for all 12 languages.
Lang. # Neurons # Trainable Param-s %
Afrikaans 3557 54.6M 0.45
Welsh 8239 126.6M 1.03
Estonian 3815 58.6M 0.48
Icelandic 5644 86.7M 0.71
Georgian 2028 31.1M 0.25
Lithuanian 3840 59.0M 0.48
Latvian 3741 57.5M 0.47
Macedonian 2403 36.9M 0.30
Maltese 3611 55.5M 0.45
Nepali 5567 85.5M 0.70
Slovak 2310 35.5M 0.29
Slovenian 2365 36.3M 0.30
Figure 11: Neuron and parameter counts for different language subnetworks, and percentages of the model being trained within Mistral-Nemo-12B.
Figure 12: Language neuron distributions for all 12 languages for Llama-3.1-8B-12B.
Figure 13: Language neuron distributions for all 12 languages for Mistral-Nemo-12B.
Appendix B Validation Losses
(a) Afrikaans validation loss curves for Llama-3.1-8B.
(b) Welsh validation loss curves for Llama-3.1-8B.
(c) Estonian validation loss curves for Llama-3.1-8B.
(d) Icelandic validation loss curves for Llama-3.1-8B.
(e) Georgian validation loss curves for Llama-3.1-8B.
(f) Lithuanian validation loss curves for Llama-3.1-8B.
(a) Latvian validation loss curves for Llama-3.1-8B.
(b) Macedonian validation loss curves for Llama-3.1-8B.
(c) Maltese validation loss curves for Llama-3.1-8B.
(d) Nepali validation loss curves for Llama-3.1-8B.
(e) Slovak validation loss curves for Llama-3.1-8B.
(f) Slovenian validation loss curves for Llama-3.1-8B.
Appendix C Downstream Results
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE afr Latn 72.11 31.22 47.67 65.00 68.89 73.11
SIB200 afr Latn 35.29 27.45 30.39 31.37 45.59 44.12
[…]
Avg. 38.01 18.47 32.73 38.40 42.03 42.99
General Performance
arc challenge 70.90 25.75 35.45 64.55 62.21 71.24
arc easy 85.79 26.14 53.33 79.12 81.75 85.26
hellaswag 40.38 14.68 26.13 35.22 36.04 38.72
piqa 71.87 39.17 50.00 73.07 72.03 71.71
winogrande 51.38 49.17 50.43 50.91 47.04 51.46
mmlu 58.07 27.63 37.30 51.60 54.74 56.43
Avg. 63.07 30.42 42.11 59.08 58.97 62.47
Table 3: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
SIB200 cym Latn 37.25 1.47 25.49 49.51 20.10 35.78
[…]
Avg. 25.43 17.95 27.00 31.33 26.61 28.65
General Performance
arc challenge 70.90 16.05 22.74 56.52 54.52 68.23
arc easy 85.79 16.84 30.70 74.56 76.14 84.04
hellaswag 40.38 12.31 25.27 29.93 30.88 40.79
piqa 71.87 3.97 49.46 58.98 66.97 68.39
winogrande 51.62 20.28 49.57 49.17 51.46 50.51
mmlu 58.07 22.27 25.67 45.79 48.14 56.69
Avg. 63.10 15.29 33.90 52.49 54.68 61.44
Table 4: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE ekk Latn 66.67 15.00 42.00 56.22 58.11 65.00
SIB200 ekk Latn 40.20 12.25 30.88 36.76 37.75 55.88
[…]
Avg. 26.57 8.60 21.55 27.81 28.99 33.30
General Performance
arc challenge 70.90 22.07 36.45 58.86 63.55 72.58
arc easy 85.79 21.40 55.09 77.89 81.58 84.74
hellaswag 40.38 13.92 22.14 36.11 37.06 41.82
piqa 71.87 40.10 54.35 59.63 72.96 75.24
winogrande 51.38 41.83 50.20 51.14 52.01 53.83
mmlu 58.07 20.64 33.90 48.66 52.84 57.35
Avg. 63.07 26.66 42.02 55.38 60.00 64.26
Table 5: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE isl Latn 62.67 21.44 41.67 50.89 51.00 66.11
SIB200 isl Latn 41.18 2.45 24.51 25.98 28.43 42.16
[…]
Avg. 26.53 11.25 21.06 23.89 24.22 29.41
General Performance
arc challenge 70.90 21.74 37.79 56.19 59.20 74.25
arc easy 85.79 23.33 55.96 74.39 76.84 85.61
hellaswag 40.38 8.29 25.24 32.65 27.75 44.67
piqa 71.87 26.55 54.79 69.48 61.48 77.31
winogrande 51.38 40.02 50.28 52.01 52.17 54.54
mmlu 58.07 17.18 35.86 46.77 45.40 58.92
Avg. 63.07 22.85 43.32 55.25 53.81 65.88
Table 6: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE kat Geor 59.67 0.00 12.67 60.33 49.56 58.89
SIB200 kat Geor 49.02 0.98 0.00 48.53 29.90 43.14
[…]
Avg. 24.16 1.88 5.64 24.04 19.27 23.05
General Performance
arc challenge 70.90 9.03 10.70 71.57 63.55 69.90
arc easy 85.79 10.18 15.44 85.26 82.98 84.56
hellaswag 40.38 1.24 6.33 42.39 33.13 41.56
piqa 71.87 12.13 8.81 73.67 64.53 71.16
winogrande 51.38 3.95 37.57 53.12 50.91 51.38
mmlu 58.07 4.51 18.42 58.00 55.26 56.69
Avg. 63.07 6.84 16.21 64.00 58.39 62.54
Table 7: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE lit Latn 66.44 34.33 46.33 53.00 64.56 67.00
SIB200 lit Latn 45.59 2.45 27.45 31.37 27.94 46.08
[…]
Avg. 29.63 11.54 21.95 26.10 28.12 31.50
General Performance
arc challenge 70.90 20.74 37.79 54.52 67.89 70.23
arc easy 85.79 27.89 59.12 74.04 80.53 85.26
hellaswag 40.38 20.49 29.02 31.45 38.14 41.37
piqa 71.87 39.28 51.63 57.56 69.15 74.21
winogrande 51.38 51.07 49.96 50.43 50.36 53.35
mmlu 58.07 27.24 36.64 45.85 52.19 57.02
Avg. 63.07 31.12 44.03 52.31 59.71 63.58
Table 8: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE lvs Latn 65.56 22.00 39.78 47.22 55.11 67.67
SIB200 lvs Latn 42.16 20.10 2.45 16.67 10.78 41.67
[…]
Avg. 29.62 13.01 19.13 24.10 25.02 32.20
General Performance
arc challenge 70.90 20.07 25.42 53.51 68.23 69.23
arc easy 85.79 27.54 36.67 70.70 81.58 84.56
hellaswag 40.38 14.48 17.34 31.23 35.03 41.73
piqa 71.87 39.50 43.69 55.93 71.82 74.59
winogrande 51.62 47.43 46.80 50.04 50.43 55.17
mmlu 58.07 24.89 30.70 45.59 53.43 59.63
Avg. 63.10 28.98 33.44 51.17 60.09 64.15
Table 9: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE mkd Cyrl 76.56 28.89 44.11 72.44 68.33 72.11
SIB200 mkd Cyrl 44.12 9.31 34.80 39.71 53.43 40.20
[…]
Avg. 33.25 12.60 24.83 33.98 36.89 34.98
General Performance
arc challenge 70.90 31.10 43.48 69.90 66.22 69.23
arc easy 85.79 39.82 64.39 84.39 82.11 85.61
hellaswag 40.38 26.41 25.06 37.91 32.57 36.43
piqa 71.87 49.67 53.26 73.72 68.44 67.36
winogrande 51.62 49.41 49.25 52.96 51.78 52.09
mmlu 58.07 30.96 39.65 56.69 54.67 57.48
Avg. 63.10 37.90 45.85 62.60 59.30 61.37
Table 10: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE mlt Latn 57.67 25.11 40.00 50.89 58.11 66.44
SIB200 mlt Latn 37.75 10.29 24.02 27.94 33.82 38.73
[…]
Avg. 30.22 14.37 30.13 27.34 34.93 39.01
General Performance
arc challenge 70.90 22.07 38.13 49.16 62.21 71.57
arc easy 85.79 21.23 56.84 72.46 82.11 83.86
hellaswag 40.38 9.70 26.72 31.45 35.07 39.00
piqa 71.87 42.11 50.87 61.97 72.85 73.34
winogrande 51.54 51.22 49.17 49.96 53.67 50.91
mmlu 58.07 26.26 32.66 45.00 52.58 57.87
Avg. 63.09 28.77 42.40 51.67 59.75 62.76
Table 11: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE npi Deva 59.67 1.22 29.44 53.89 55.78 57.44
SIB200 npi Deva 36.27 4.41 4.41 27.45 33.33 57.84
[…]
Avg. 26.89 4.74 12.93 26.54 27.45 32.65
General Performance
arc challenge 70.90 12.71 35.45 65.89 67.56 71.57
arc easy 85.79 19.47 49.12 81.75 82.98 84.39
hellaswag 40.38 16.38 14.32 37.46 36.68 36.31
piqa 71.87 36.83 51.25 62.35 73.83 71.11
winogrande 51.38 13.58 47.04 50.12 54.06 50.99
mmlu 58.07 17.50 33.18 54.08 54.08 57.74
Avg. 63.07 19.41 38.39 58.61 61.53 62.02
Table 12: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE slk Latn 73.67 27.44 39.78 61.11 65.56 72.89
SIB200 slk Latn 40.20 20.59 39.22 44.61 33.82 46.08
[…]
Avg. 33.48 16.86 26.49 32.76 32.44 35.56
General Performance
arc challenge 70.90 21.74 36.45 61.20 62.21 71.57
arc easy 85.79 23.86 51.05 78.25 78.60 85.44
hellaswag 40.38 15.04 27.87 40.04 36.74 43.68
piqa 71.87 43.25 54.30 59.25 70.29 71.76
winogrande 51.54 38.91 50.43 51.07 50.91 51.70
mmlu 58.07 21.36 35.99 52.25 51.53 57.67
Avg. 63.09 27.36 42.68 57.01 58.38 63.64
Table 13: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Full MLP LoRA Random Target
Monolingual Capabilities
[…]
BELEBELE slv Latn 71.00 26.78 38.44 51.67 63.00 68.56
SIB200 slv Latn 41.67 23.04 16.18 20.59 53.92 47.55
[…]
Avg. 30.88 11.32 18.31 24.66 32.75 32.90
General Performance
arc challenge 70.90 21.74 34.78 53.51 67.56 67.89
arc easy 85.79 28.60 36.49 71.23 82.46 83.51
hellaswag 40.38 8.79 24.92 34.08 39.72 36.48
piqa 71.87 45.65 56.15 70.13 72.03 75.08
winogrande 51.46 49.88 50.67 49.64 50.43 52.25
mmlu 58.07 27.43 29.13 43.31 55.06 56.69
Avg. 63.08 30.35 38.69 53.65 61.21 61.98
Table 14: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Llama-3.1-8B. Best performing configurations are highlighted in dark green, second-best in light green.
C.1 Mistral-Nemo-12B
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE afr Latn 79.67 75.56 79.44
SIB200 afr Latn 31.86 43.14 31.86
[…]
Avg. 35.11 39.93 38.57
General Performance
arc challenge 80.27 72.24 75.59
arc easy 89.82 83.33 87.72
hellaswag 40.86 38.31 42.32
piqa 80.03 77.15 77.20
winogrande 54.22 53.28 52.17
mmlu 61.66 55.85 61.33
Avg. 67.81 63.36 66.06
Table 15: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
SIB200 cym Latn 35.78 31.37 40.20
[…]
Avg. 21.41 30.09 31.77
General Performance
arc challenge 80.27 63.88 71.57
arc easy 89.82 78.77 85.09
hellaswag 40.86 30.91 35.44
piqa 80.03 61.15 65.61
winogrande 54.22 51.07 51.93
mmlu 61.66 50.10 55.58
Avg. 67.81 55.98 60.87
Table 16: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE ekk Latn 73.00 62.44 71.00
SIB200 ekk Latn 32.35 29.90 36.76
[…]
Avg. 25.18 26.53 28.37
General Performance
arc challenge 80.27 71.91 74.92
arc easy 89.82 83.68 88.42
hellaswag 40.86 32.99 40.67
piqa 80.03 68.61 77.75
winogrande 54.22 52.09 53.35
mmlu 61.66 54.93 59.83
Avg. 67.81 60.70 65.82
Table 17: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE isl Latn 67.00 60.00 70.67
SIB200 isl Latn 34.80 25.98 38.24
[…]
Avg. 24.75 19.25 29.14
General Performance
arc challenge 80.27 64.55 76.59
arc easy 89.82 83.16 88.95
hellaswag 40.86 31.59 41.73
piqa 80.03 60.72 74.54
winogrande 54.22 51.70 53.67
mmlu 61.66 51.99 60.94
Avg. 67.81 57.28 66.07
Table 18: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE kat Geor 71.89 68.56 72.33
SIB200 kat Geor 30.39 37.25 32.35
[…]
Avg. 21.99 24.50 23.00
General Performance
arc challenge 80.27 69.23 77.26
arc easy 89.82 82.81 87.37
hellaswag 40.86 41.17 37.87
piqa 80.03 70.08 72.47
winogrande 54.22 52.33 53.99
mmlu 61.66 56.63 60.48
Avg. 67.81 62.04 64.91
Table 19: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lit Latn 73.44 64.11 71.56
SIB200 lit Latn 33.82 25.98 34.80
[…]
Avg. 26.55 25.46 28.61
General Performance
arc challenge 80.27 68.56 76.59
arc easy 89.82 86.32 88.77
hellaswag 40.85 36.47 41.65
piqa 80.03 67.25 72.20
winogrande 54.22 51.14 53.83
mmlu 61.66 52.25 60.16
Avg. 67.81 60.33 65.53
Table 20: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lvs Latn 72.89 60.44 66.00
SIB200 lvs Latn 34.80 46.08 30.88
[…]
Avg. 27.55 31.21 27.39
General Performance
arc challenge 80.27 64.21 72.91
arc easy 89.82 83.68 87.54
hellaswag 40.86 34.50 35.25
piqa 80.03 66.32 62.84
winogrande 54.22 51.22 51.93
mmlu 61.66 53.89 58.79
Avg. 67.81 58.97 61.54
Table 21: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mkd Cyrl 76.11 73.56 76.67
SIB200 mkd Cyrl 31.37 31.86 44.12
[…]
Avg. 29.82 32.59 32.47
General Performance
arc challenge 80.27 73.58 75.92
arc easy 89.82 87.02 87.89
hellaswag 40.86 37.81 44.30
piqa 80.03 74.97 76.17
winogrande 54.22 51.93 54.06
mmlu 61.66 57.28 59.57
Avg. 67.81 63.77 66.32
Table 22: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mlt Latn 41.33 64.67 72.44
SIB200 mlt Latn 30.88 44.61 33.33
[…]
Avg. 18.89 36.86 36.60
General Performance
arc challenge 80.27 68.23 74.92
arc easy 89.82 80.88 86.67
hellaswag 40.86 35.46 39.45
piqa 80.03 72.74 65.23
winogrande 54.22 51.54 52.57
mmlu 61.66 52.45 59.24
Avg. 67.81 60.22 63.01
Table 23: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE npi Deva 64.11 51.56 60.67
SIB200 npi Deva 34.31 12.25 22.55
[…]
Avg. 25.56 21.35 23.72
General Performance
arc challenge 80.27 51.51 72.58
arc easy 89.82 66.84 87.19
hellaswag 40.86 29.48 35.37
piqa 80.03 63.38 76.01
winogrande 54.22 52.49 53.12
mmlu 61.66 46.18 55.45
Avg. 67.81 51.65 63.29
Table 24: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE slk Latn 79.78 71.89 77.33
SIB200 slk Latn 34.80 44.12 28.43
[…]
Avg. 30.04 34.08 32.01
General Performance
arc challenge 80.27 70.90 75.59
arc easy 89.82 85.09 88.77
hellaswag 40.86 36.49 40.12
piqa 80.03 70.95 78.62
winogrande 54.22 51.62 55.49
mmlu 61.66 53.36 60.42
Avg. 67.81 61.40 66.50
Table 25: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak. Best performing configurations are highlighted in dark green, second-best in light green.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE slv Latn 77.56 68.33 71.89
SIB200 slv Latn 32.84 32.35 26.96
[…]
Avg. 27.36 26.64 24.93
General Performance
arc challenge 80.27 68.56 72.91
arc easy 89.82 83.51 84.74
hellaswag 40.86 34.75 37.04
piqa 80.03 64.58 75.24
winogrande 54.22 53.04 54.14
mmlu 61.66 55.13 57.41
Avg. 67.81 59.93 63.58
Table 26: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian. Best performing configurations are highlighted in dark green, second-best in light green.
C.2 Generation Prompts
Task Prompts
We use standardized prompt templates for each task category to ensure consistency across models and datasets. Below we show the representative formats. Placeholders such as {source_text}, {passage}, or {question} are replaced with the actual content at runtime.
Translation.
Translate this sentence into {target_name}: {source_text} Translation:
Classification.
Classify the following text into one of these topics:
A) Science/Technology
B) Travel
C) Politics
D) Sports
E) Health
F) Entertainment
G) Geography
Text: {text}
Answer:
Comprehension.
Read the following passage and pick the right answer.
Passage: {passage}
Question: {question}
{options}
Answer:
General Capabilities.
We adopt dataset-specific templates for MMLU and commonsense reasoning benchmarks:
  • MMLU:
    Answer the following multiple choice question:
    {question}
    {options}
    Answer:
  • HellaSwag:
    Complete the following scenario by choosing the most likely continuation:
    {ctx}
    {options}
    Answer:
  • PIQA:
    Choose the most appropriate solution for the following goal:
    Goal: {goal}
    A) {sol1}
    B) {sol2}
    Answer:
  • Winogrande:
    Complete the sentence by choosing the correct option:
    {sentence}
    A) {option1}
    B) {option2}
    Answer:
  • ARC (Easy/Challenge):
    Answer the following question:
    {question}
    {options}
    Answer:
C.3 Generation Settings
The generation is done using the following parameters:
SamplingParams(    temperature=0,    repetition_penalty=1.1,    max_tokens=128,    top_p=1.0,    top_k=-1,    stop_token_ids=[eos_token_id] if eos_token_id is not None else [],    stop=[
[…]
    "Question:", "Q:",
    "Answer:", "A:",
    "Translation:", "The translation", "Translate:",
    "Summary:", "The summary", "Summarize:",
    "Passage:", "The passage", "Text:",
    "Options:", "The options", "The choices", "Choices:",
    "Context:", "The context", "The problem", "Problem:",
[…]
    "</s>", "<|endoftext|>",
    "Human:", "Assistant:"
   skip_special_tokens=True
Appendix D Cross-lingual Alignment
ar de en es fr it ja ru zh
Baseline 85.55 92.88 90.63 91.47 91.21 91.45 86.55 88.21 87.46
Full 86.73 91.90 94.65 91.14 91.48 89.98 86.10 87.36 88.83
MLP 84.22 90.39 94.05 89.85 90.39 89.62 85.06 87.48 88.19
Adapter 84.79 95.11 92.10 93.23 93.21 93.07 86.58 88.59 87.47
Random 86.76 95.06 92.94 92.84 92.95 92.66 88.15 90.18 89.31
Target 87.75 96.10 95.03 94.24 94.52 93.89 90.24 91.25 91.46
Table 27: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Afrikaans) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 83.22 87.82 85.92 87.72 87.58 88.11 83.09 84.76 83.24
Full 88.11 90.77 92.58 91.70 91.88 92.55 87.44 87.37 89.85
MLP 84.88 90.89 91.98 89.65 90.63 90.65 85.80 87.55 88.96
Adapter 88.13 93.40 90.40 93.30 93.29 94.08 86.11 89.41 86.25
Random 89.94 92.71 89.41 92.48 92.39 92.13 86.22 89.49 86.88
Target 94.35 95.49 93.20 94.90 95.57 95.62 91.70 92.71 91.57
Table 28: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Welsh) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 85.45 91.29 89.18 90.46 90.19 90.67 86.09 88.40 86.75
Full 83.70 91.91 93.61 91.33 91.97 93.14 85.91 87.06 89.27
MLP 80.59 90.15 92.08 87.14 88.69 89.03 82.08 86.85 85.91
Adapter 86.19 93.87 90.52 92.56 92.51 92.83 86.99 89.58 87.47
Random 84.61 93.27 89.25 91.13 91.01 91.38 84.89 89.54 85.50
Target 90.07 95.43 93.87 94.13 93.91 93.73 90.77 92.25 91.84
Table 29: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Estonian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 82.82 88.54 85.85 87.65 87.45 87.89 83.22 85.24 83.55
Full 89.21 90.85 92.80 90.94 90.99 88.86 88.24 88.52 90.31
MLP 80.94 90.20 90.23 88.49 88.76 89.58 78.67 84.65 82.93
Adapter 85.49 93.79 89.50 92.61 92.24 92.57 85.22 88.89 86.21
Random 87.74 91.14 89.89 91.65 90.48 90.17 85.29 91.03 87.21
Target 91.14 93.44 93.11 94.09 94.10 94.25 90.11 91.98 90.65
Table 30: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Icelandic) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 66.64 68.23 61.90 66.87 67.40 68.37 66.58 70.69 64.28
Full 64.74 64.17 57.36 64.37 63.03 64.32 63.36 66.32 61.19
MLP 67.84 69.05 61.96 67.05 67.74 68.95 67.34 69.48 65.94
Adapter 65.52 67.76 61.55 66.33 66.81 67.67 65.41 70.47 63.65
Random 67.02 69.26 62.24 67.80 68.14 69.37 66.86 71.79 65.19
Target 69.54 70.67 64.23 69.01 69.66 70.64 69.20 73.41 67.32
Table 31: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Georgian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 83.79 89.55 86.95 89.12 88.88 89.56 84.84 87.59 85.20
Full 86.30 92.68 92.10 91.05 91.04 91.40 88.21 87.89 89.71
MLP 82.28 91.34 91.10 89.48 89.88 90.35 85.49 86.40 87.62
Adapter 83.45 91.75 88.31 91.17 90.83 91.67 84.16 88.73 84.43
Random 83.26 91.59 88.84 91.37 90.73 91.36 84.83 88.83 85.68
Target 86.99 93.23 91.85 92.78 92.50 92.47 88.56 91.26 90.07
Table 32: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Lithuanian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 82.73 88.36 85.37 87.88 87.58 88.25 83.37 86.37 83.76
Full 86.70 89.06 90.55 90.20 89.57 88.98 87.45 84.64 89.21
MLP 80.97 88.03 88.93 86.05 86.14 86.56 82.48 84.67 86.19
Adapter 84.62 92.14 88.45 91.85 91.36 92.08 84.51 89.09 86.13
Random 82.07 91.14 87.65 90.65 90.18 90.39 83.74 87.63 84.94
Target 86.60 93.48 91.73 93.19 93.04 93.14 88.61 90.79 89.61
Table 33: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Latvian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 87.36 87.33 86.16 87.00 86.85 87.38 86.00 92.75 86.12
Full 85.37 89.55 90.10 86.95 89.35 87.79 84.21 92.15 87.69
MLP 84.71 88.09 90.33 87.73 87.95 88.06 86.25 90.71 88.38
Adapter 87.53 86.66 85.75 86.49 86.15 86.97 84.81 97.54 85.33
Random 94.88 89.60 87.52 89.51 89.00 90.22 88.70 98.18 88.11
Target 92.74 92.92 92.01 92.99 92.63 93.00 91.52 98.72 92.00
Table 34: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Macedonian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 84.92 87.62 85.09 87.37 87.32 89.50 82.94 84.95 83.17
Full 86.03 90.98 91.14 89.55 89.71 90.60 85.43 88.60 85.49
MLP 86.29 90.72 90.80 90.55 89.50 90.07 86.65 88.56 87.35
Adapter 88.20 93.48 89.95 93.85 93.66 96.17 85.00 88.92 85.74
Random 89.54 91.84 90.38 91.49 91.17 93.10 86.12 89.70 86.50
Target 92.23 93.53 92.42 92.79 92.63 94.90 89.06 90.80 89.52
Table 35: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Maltese) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 85.43 84.39 82.60 83.91 83.85 84.81 85.77 86.33 84.49
Full 88.65 88.61 89.58 89.29 89.63 89.65 88.12 89.33 89.39
MLP 83.47 83.90 86.27 84.17 84.67 83.77 85.09 85.00 86.68
Adapter 89.31 86.91 84.48 86.07 85.95 86.98 88.93 89.12 86.84
Random 92.35 87.82 84.56 87.06 86.96 87.95 90.76 91.25 88.24
Target 94.93 92.81 89.43 91.40 91.43 92.10 94.87 94.62 92.49
Table 36: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Nepali) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 87.04 91.90 90.14 91.61 91.24 91.61 87.26 90.57 87.85
Full 87.16 92.87 93.77 91.80 91.93 91.83 88.86 88.68 89.80
MLP 84.25 91.34 93.15 90.85 90.58 90.91 84.58 88.15 87.52
Adapter 85.71 92.33 90.44 92.23 91.83 92.12 86.20 90.18 87.07
Random 87.06 92.57 91.00 92.11 91.82 91.69 87.97 91.88 88.57
Target 87.56 93.88 93.35 93.52 93.25 93.33 89.86 92.60 90.45
Table 37: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Slovak) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 85.88 91.34 89.69 91.13 90.40 91.31 86.13 89.15 86.98
Full 86.41 89.81 93.60 90.56 89.94 90.11 86.23 89.84 88.16
MLP 84.26 89.02 92.18 89.38 89.64 89.61 84.05 87.33 87.56
Adapter 86.40 92.67 91.10 92.64 92.03 92.60 86.11 90.97 87.25
Random 84.48 92.45 90.52 91.82 91.10 92.14 84.45 91.15 85.49
Target 88.57 94.18 93.79 94.25 93.47 94.17 88.72 93.27 90.21
Table 38: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Slovenian) and other languages for Llama-3.1-8B.
Appendix E Computational Budget
All experiments were conducted on a single NVIDIA H100 GPU, except for the full fine-tuning setup, which required two H100 GPUs due to its higher memory and compute demands. Each model was trained for approximately 100 million tokens, corresponding to about 6-8 GPU hours per run for parameter-efficient methods and roughly 10 GPU hours for full fine-tuning.
Appendix F Licenses and Attribution
All experiments were conducted using publicly available models and datasets released under permissive or research-oriented licenses. The Llama-3.1-8B model is distributed under the Meta Llama 3 Community License Agreement, which permits research use and redistribution with attribution. Mistral-Nemo-12B is released under the Apache License 2.0, allowing open use and modification. Language-specific token data were obtained from the GlotCC corpus (CC0 for metadata, Common Crawl terms for text content). Evaluation datasets including FLORES, BELEBELE, SIB200, and OPUS were used in accordance with their respective research licenses, and English benchmarks (ARC, Hellaswag, PIQA, Winogrande, MMLU) were accessed under standard academic usage terms.