### v2

# Sparse Subnetwork Enhancement for Underrepresented Languages in Large Language Models

Abstract
Large language models (LLMs) exhibit substantial performance disparities across languages, particularly between high- and low-resource settings. We propose a framework for improving performance in underrepresented languages while preserving general-purpose capabilities via targeted fine-tuning of sparse, language-associated subnetworks. Our approach identifies language-relevant neurons using Language Activation Probability Entropy (LAPE), an information-theoretic metric that reliably captures language-specific activation patterns, and fine-tunes only the corresponding weights. Experiments on Llama-3.1-8B, Mistral-Nemo-12B, and Aya-Expanse-8B across 12 mid- and low-resource languages show that our method consistently outperforms full fine-tuning, FFN-only fine-tuning, LoRA, IA3, and random-subset baselines while updating only 0.2–1% of model parameters. We further show that sparse, neuron-targeted fine-tuning can inject new language capabilities without catastrophic forgetting, with potential applicability to other model capabilities. Mechanistic analyses of weight updates and internal representations reveal asymmetric roles of FFN projections in language adaptation and improved cross-lingual alignment. Finally, we release language neuron sets for over 100 languages together with our adaptation pipeline, enabling a cost-effective path for extending LLMs to underrepresented languages.
Machine Learning, ICML
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks and languages, yet their performance remains unevenly distributed, with substantial gaps between high-resource and low-resource languages (Joshi et al., 2020; Huang et al., 2023). While scaling training data and model parameters has improved multilingual performance, this approach is often impractical for resource-constrained scenarios involving hundreds of languages (Robinson et al., 2023; Lai et al., 2024).
Multi-HeadAttentionFeed-Forward NetworkTrainedFrozen
Figure 1: Sparse subnetwork adaptation. Attention blocks remain frozen while LAPE-identified language-specific neurons and their weights within FFNs are fine-tuned.
Recent advances in mechanistic interpretability have revealed that language models develop specialized internal structures for different languages, with distinct neurons and attention patterns activated for specific linguistic contexts (Tang et al., 2024; Zhao et al., 2024; Liu et al., 2024; Kojima et al., 2024). This observation suggests a promising alternative to full model retraining: identifying and selectively enhancing language-specific components while preserving general-purpose capabilities.
Current approaches to multilingual adaptation face a fundamental trade-off between target-language performance and retention of general capabilities. Full fine-tuning often leads to catastrophic forgetting (French, 1999; Kirkpatrick et al., 2017), while parameter-efficient methods (PEFT) (Houlsby et al., 2019; Hu et al., 2022; Vykopal et al., 2025) have emerged as practical alternatives; however, they operate without explicit consideration of how knowledge is organized within the model, which may lead to the indiscriminate modification of parameters encoding different functions. Neuron-level approaches have shown promise in this regard but often underperform PEFT methods. Moreover, they remain largely unexplored for systematic language enhancement: prior work has focused on task-specific adaptation, such as machine translation (Zhu et al., 2024; Xu et al., 2025), or employed neuron identification primarily for interpretability rather than targeted fine-tuning (Tang et al., 2024; Kojima et al., 2024; Wang et al., 2025). A key question remains open: can we inject new language capabilities into LLMs by directly fine-tuning only the sparse subset of neurons that encode language-specific processing?
In this work, we provide evidence that this is indeed feasible. We present a framework for enhancing monolingual capabilities of LLMs through targeted fine-tuning of language-specific subnetworks within feed-forward network (FFN) components. We focus on FFNs, which prior work has shown to be the primary site of linguistic knowledge in transformers (Geva et al., 2021; Wang et al., 2022). Our approach uses Language Activation Probability Entropy (LAPE) (Tang et al., 2024), an information-theoretic metric that quantifies how selectively neurons activate for specific languages, to identify neurons exhibiting strong language-specific patterns within Llama-3.1-8B (Grattafiori et al., 2024), Mistral-Nemo-12B (Mistral AI Team, 2024), and Aya-Expanse-8B (Dang et al., 2024). We then fine-tune only the weights associated with these neurons using target-language data, updating just 0.2–1% of model parameters.
Our central finding is that sparse, targeted fine-tuning of language-specific neurons can systematically inject new language capabilities without catastrophic forgetting while outperforming common PEFT methods, a result that, to our knowledge, has not been demonstrated at this scale. Beyond performance gains, we provide mechanistic insights into how language adaptation occurs: weight-change analysis reveals asymmetric roles of FFN projection matrices, with down-projections in later layers showing substantially larger updates, suggesting they serve as the primary site of linguistic re-wiring during adaptation.
We make the following contributions:
  • We introduce a systematic framework for enhancing underrepresented languages through sparse subnetwork fine-tuning, demonstrating that language capabilities can be injected by updating less than 1% of parameters.
  • We provide empirical validation across 12 mid- and low-resource languages on three models, showing (i) consistent absolute improvements over full fine-tuning, FFN-only fine-tuning, LoRA, IA3 (Liu et al., 2022), and random subset baselines and (ii) full preservation of general capabilities, which is otherwise only exhibited by IA3.
  • We offer mechanistic insights into language adaptation: cross-lingual alignment improves systematically, and weight changes concentrate in later-layer down-projections, revealing their central role in integrating new language-specific knowledge.
  • We release language-specific neuron identifications for over 100 languages alongside our adaptation pipeline, providing practical resources for the community. [^1] Code, neuron indices, and adapted models are available on GitHub and Hugging Face.
This work represents a step toward democratizing high-quality language model capabilities. By demonstrating that principled subnetwork identification enables effective adaptation with minimal computational resources, we provide a practical framework and a foundation for unlocking future research on LLMs for underrepresented languages.
2 Related Work
Prior work on adapting multilingual language models to underrepresented languages can be grouped into three main directions: full fine-tuning, adapter-based approaches, and more recent neuron- or head-level methods.
2.1 Full Fine-Tuning Approaches
Full fine-tuning is the most direct strategy for language adaptation. Post-training on related languages has been used to mitigate overfitting (Neubig & Hu, 2018), while domain-adaptive fine-tuning improved contextualized models on specialized corpora (Han & Eisenstein, 2019). Language-specific fine-tuning on monolingual corpora (Gururangan et al., 2020; Chau et al., 2020), transliteration-based adaptation (Muller et al., 2021), and even small corpora such as Bible translations (Ebrahimi & Kann, 2021) have boosted performance on tagging, parsing, and NER. However, full fine-tuning is costly, data-hungry, and prone to catastrophic forgetting (French, 1999), motivating parameter-efficient alternatives.
2.2 Parameter-Efficient Approaches
Adapters (Houlsby et al., 2019) introduce small trainable modules into frozen backbones, substantially reducing computational cost (Strubell et al., 2019) and data requirements while mitigating catastrophic forgetting (French, 1999). Prominent multilingual extensions such as MAD-X (Pfeiffer et al., 2020), UDapter (Üstün et al., 2020), and MAD-G (Ansell et al., 2021) incorporate language- and task-specific modularity, typological signals, or hierarchical structure, while other work explores language balancing (Lee et al., 2022a; Parović et al., 2022) and inference-aware training (Parovic et al., 2023). Recent studies further show that adapter tuning can outperform continued pre-training for language adaptation (Yong et al., 2023; Gurgurov et al., 2025b). Beyond adapters, several parameter-efficient alternatives have been proposed. IA3 (Liu et al., 2022) scales activations using learned vectors, Compacter (Karimi Mahabadi et al., 2021) employs low-rank hypercomplex adapters, and AdaLoRA (Zhang et al., 2023) dynamically allocates rank budgets during training. While effective for task adaptation, these approaches introduce auxiliary parameters and do not explicitly leverage or modify the model’s internal multilingual structure.
2.3 Neuron-Level Approaches
Recent work has explored fine-grained adaptation by targeting individual model components. Early approaches used gradient-based criteria, such as magnitude pruning (Han et al., 2015), movement pruning (Sanh et al., 2020), Taylor expansion scores (Molchanov et al., 2017), to identify important parameters; though these were developed for older CNN and encoder-only architectures. For modern LLMs, SLAM (Fan et al., 2025) selects layers with high multilingual activation variance and fine-tunes their FFN sublayers, but operates at layer granularity. Zhu et al. (2024) adapt Taylor scores to identify language-aware components for machine translation, but require gradient computation on parallel data, limiting applicability to low-resource settings. Zhao et al. (2024) use ablation studies to identify neurons and fine-tune $\sim$0.1% of parameters for four high-resource languages, but report rapid saturation with additional data due to overfitting to the training source. Xu et al. (2025) propose NeFT, identifying task-sensitive neurons by comparing pre- and post-fine-tuning representations, though this requires a full fine-tuning pass before neuron selection. Mondal et al. (2025) apply LAPE for neuron identification but observe limited gains; notably, their study focuses on three high-resource languages and adapts LoRA modules atop FFNs rather than directly fine-tuning neuron weights. previous work has not succeeded in using these structures reliably
2.4 Positioning of Our Work
Despite growing evidence that language-specific structures exist inside LLMs, previous work has not succeeded in using these structures reliably to add new language capabilities at scale without degrading existing ones. Existing neuron identification methods rely on task-specific signals: Taylor scores require gradient computation on parallel data (Zhu et al., 2024), NeFT requires a fine-tuning pass to identify task-sensitive neurons (Xu et al., 2025), and ablation-based methods (Zhao et al., 2024) are computationally expensive at scale. In contrast, LAPE (Tang et al., 2024) offers a task-agnostic, information-theoretic measure of language selectivity derived directly from activation statistics and requires only monolingual text. Prior work has shown that LAPE enables robust neuron identification and target manipulation across languages (Tang et al., 2024; Gurgurov et al., 2025a). Building on this foundation, we establish sparse subnetwork fine-tuning as a viable mechanism for capability-level language adaptation, rather than task-specific adjustment, while explicitly maintaining general model competence. We demonstrate that this approach holds across models, languages, and benchmarks.
3 Methodology
Our approach consists of two parts: (i) identifying language-sensitive neurons, and (ii) selectively fine-tuning the corresponding subnetwork (Figure 1). This enables adaptation to underrepresented languages while minimizing updates to unrelated parameters.
3.1 Neuron Identification
We leverage Language Activation Probability Entropy (LAPE) (Tang et al., 2024) to identify neurons exhibiting language-specific behavior. As discussed in Section 2, LAPE is task-agnostic and requires only monolingual text, making it well-suited to low-resource settings. We focus on neurons within FFN components, as prior work has established that feed-forward layers serve as the primary focus of factual and linguistic knowledge storage (Geva et al., 2021; Dai et al., 2022; Meng et al., 2022), whereas attention heads primarily mediate information routing and contextual dependencies (Voita et al., 2019; Clark et al., 2019).
We compute LAPE over a reference set of $K=12$ typologically diverse underrepresented languages (Joshi et al., 2020): Maltese, Afrikaans, Icelandic, Welsh, Macedonian, Latvian, Lithuanian, Slovenian, Slovak, Estonian, Georgian, and Nepali.
For each neuron $i$ in layer $\ell$, we estimate the probability of activation under inputs from language $k$:
$p^{k}_{\ell,i}=\mathbb{E}_{x\sim\mathcal{D}_{k}}\Big[\mathbf{1}\!\left(\sigma(W_{\text{gate}}^{(\ell)}h^{(\ell)}(x))_{i}>0\right)\Big],$ (1)
where $\mathcal{D}_{k}$ is a monolingual corpus for language $k$, $h^{(\ell)}(x)\in\mathbb{R}^{d}$ is the hidden state input to layer $\ell$, $W_{\text{gate}}^{(\ell)}\in\mathbb{R}^{d_{\text{ff}}\times d}$ is the gate projection matrix, $\sigma(\cdot)$ is the SiLU activation function, and $\mathbf{1}(\cdot)$ is the indicator function.
The probability vector across all languages $k\in\{1,\dots,K\}$ is normalized to obtain a distribution:
$\tilde{p}_{\ell,i}^{k}=\frac{p^{k}_{\ell,i}}{\sum_{k^{\prime}=1}^{K}p^{k^{\prime}}_{\ell,i}}.$ (2)
The LAPE score is then computed as the Shannon entropy (Shannon & Weaver, 1998) of this distribution:
$\text{LAPE}_{\ell,i}=-\sum_{k=1}^{K}\tilde{p}_{\ell,i}^{k}\log\tilde{p}_{\ell,i}^{k}.$ (3)
Low LAPE scores indicate language-specific neurons, as their activations concentrate on a small subset of languages. We select neurons in the lowest $\rho$-th percentile of LAPE scores ($\rho=5\%$), subject to an activity threshold $\tau_{\text{act|}|}$, defined as the 95th percentile of activation probabilities across all neurons and languages. This threshold is applied in two ways: (1) globally, we require $\max_{k}p^{k}_{\ell,i}\geq\tau_{\text{act}}$ to discard uniformly inactive neurons that achieve low entropy trivially; (2) per-language, we assign a neuron to the subnetwork $\mathcal{S}^{k}$ only if $p^{k}_{\ell,i}\geq\tau_{\text{act|}|}$.
Since assignment is based on exceeding the activation threshold rather than exclusive assignment to the maximally-activating language, neurons with high activation across multiple languages may belong to multiple subnetworks. This reflects the linguistic reality that some neurons encode features shared across related languages.
3.2 Subnetwork Fine-Tuning
Given the subnetwork $\mathcal{S}^{k}$ for language $k$, we restrict parameter updates to the weights associated with the selected neurons. We consider transformer architectures with GLU-based feed-forward blocks (Shazeer, 2020), which contain three weight matrices per layer $\ell$:
  • Gate projection: $W_{\text{gate}}^{(\ell)}\in\mathbb{R}^{d_{\text{ff}}\times d}$
  • Up projection: $W_{\text{up}}^{(\ell)}\in\mathbb{R}^{d_{\text{ff}}\times d}$
  • Down projection: $W_{\text{down}}^{(\ell)}\in\mathbb{R}^{d\times d_{\text{ff}}}$
where $d$ is the hidden dimension and $d_{\text{ff}}$ is the intermediate dimension. The FFN computation for input $\mathbf{h}\in\mathbb{R}^{d}$ is given by:
$\text{FFN}(\mathbf{h})=W_{\text{down}}\cdot\Big(\sigma(W_{\text{gate}}\mathbf{h})\odot(W_{\text{up}}\mathbf{h})\Big),$ (4)
where $\sigma(\cdot)$ denotes the SiLU activation function and $\odot$ represents element-wise multiplication.
For each neuron $i$ in the subnetwork $\mathcal{S}^{k,(\ell)}$ at layer $\ell$, we define the associated trainable parameters as:
$\theta_{i}^{(\ell)}=\left\{W_{\text{gate}}^{(\ell)}[i,:],\;W_{\text{up}}^{(\ell)}[i,:],\;W_{\text{down}}^{(\ell)}[:,i]\right\},$ (5)
where $W[i,:]$ denotes the $i$-th row and $W[:,i]$ denotes the $i$-th column. The complete set of trainable parameters for language $k$ is then:
$\theta_{\mathcal{S}^{k}}=\bigcup_{\ell=1}^{L}\bigcup_{i\in\mathcal{S}^{k,(\ell)}}\theta_{i}^{(\ell)}.$ (6)
Fine-tuning is performed by minimizing the language modeling loss on target-language data $\mathcal{D}_{\text{target}}^{k}$:
$\min_{\theta_{\mathcal{S}^{k}}}\;\mathbb{E}_{x\sim\mathcal{D}_{\text{target}}^{k}}\left[-\sum_{t=1}^{|x|}\log P(x_{t}\mid x_{<t};\theta)\right],$ (7)
while keeping all remaining parameters $\theta\setminus\theta_{\mathcal{S}^{k}}$ frozen. This selective optimization adapts only the language-relevant subnetwork while preserving the model’s general capabilities.
4 Experimental Setup
Below, we describe the details of our training setup, the evaluation tasks chosen to assess model performance, and the baselines against which we compare.
4.1 Training Setup
We experiment with Llama-3.1-8B, Mistral-Nemo-12B and Aya-Expanse-8B, using GlotCC (Kargaran et al., 2024) as target-language training data. Each subnetwork is trained on up to 100M tokens by default, with additional experiments varying the size from 50M to 200M tokens for a subset of languages. We focus on 12 underrepresented languages (Joshi et al., 2020), reported in Table 1.
Subnetworks are identified for each language using 100MB of the target data, with thresholds $\rho=5\%$ and activity thresholds of 0.95. Each subnetwork is trained for one epoch (batch size 2) using AdamW (Loshchilov & Hutter, 2019) with a learning rate of $1\times 10^{-4}$, weight decay 0.01, and gradient clipping (Mikolov et al., 2012). We validate every 500 steps and select the best checkpoint via validation loss on the corresponding Flores-200 subset (Costa-Jussà et al., 2022).
Lang. # Neurons # Param-s % Model
Afrikaans 2581 31.7M 0.39
Welsh 6020 74.0M 0.92
Estonian 3204 39.4M 0.49
Icelandic 5033 61.8M 0.77
Georgian 1150 14.1M 0.18
Lithuanian 2978 36.6M 0.46
Latvian 3212 39.5M 0.49
Macedonian 1743 21.4M 0.27
Maltese 3015 37.0M 0.46
Nepali 4625 56.8M 0.71
Slovak 1940 23.8M 0.30
Slovenian 2046 25.1M 0.31
Table 1: Language subnetwork statistics for Llama-3.1-8B: neuron counts, trainable parameters, and percentage of total model parameters.
4.2 Evaluation Tasks
We evaluate our method along two dimensions:
Figure 2: Neuron distributions for Llama-3.1-8B.
Target-language performance. We assess translation quality on Flores-200 (“valtest”) and OPUS-100 (“test”) via Sentence BLEU (Post, 2018), while comprehension and classification are evaluated on Belebele (Bandarkar et al., 2024) and SIB-200 (Adelani et al., 2024) using accuracy.
General capabilities. We measure the effect of adaptation on general capabilities via accuracy on several benchmarks: MMLU (Hendrycks et al., 2021), Hellaswag (Zellers et al., 2019), PIQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2019), ARC (Clark et al., 2018).
For all tasks, we use deterministic generation with zero temperature to eliminate variance introduced by stochastic decoding. The prompts and hyperparameters specific to each evaluation are listed in Appendix E.
4.3 Baselines
We compare our sparse fine-tuning approach against five baselines:
  • (i) Full FT, full model fine-tuning ,
  • (ii) FFN FT, fine-tuning only the FFN components,
  • (iii) Rand. N-s, fine-tuning a random set of FFN neurons, uniformly sampled across layers, matching the size of the LAPE-selected subnetwork,
  • (iv) LoRA (Hu et al., 2022) applied to gate_proj, up_proj, and down_proj, configured to match our average parameter count (rank 62, alpha 32, yielding 57M trainable parameters, 0.7%), and
  • (v) IA3 (Liu et al., 2022) applied to the standard target modules (k_proj, v_proj, down_proj), which learns rescaling vectors rather than low-rank updates, providing approximately 0.5M trainable parameters (0.006%).
We include both LoRA and IA3 as parameter-efficient baselines: LoRA remains the most widely adopted PEFT method, while IA3 represents an extreme point on the parameter-efficiency spectrum, achieving competitive performance with orders of magnitude fewer parameters.
For full-model and FFN-only fine-tuning, we use a learning rate of $1\times 10^{-5}$ to avoid instability from the larger parameter count. For LoRA, IA3, and the random-subset baseline, we use identical hyperparameters to our method ($1\times 10^{-4}$). Consistent with Schulman & Lab (2025), we find that PEFT methods require learning rates approximately $10\times$ larger than full fine-tuning; using lower rates results in undertraining under our fixed budget.
5 Experimental Results
In this section, we present our experimental findings. We first analyze language-specific subnetwork identification, examining the extent to which distinct subnetworks emerge across languages. We then discuss subnetwork fine-tuning and its impact on performance relative to baselines.
Metric Base Full FT FFN FT LoRA IA3 Rand. N-s Lang. N-s
% Param-s tuned 100% 70% 0.7% 0.006% 0.2–1% 0.2–1%
Monolingual Capabilities
FLORES tg→en 28.82 11.44 25.06 28.00 29.61 30.95 31.37
FLORES en→tg 11.62 9.29 13.70 15.60 12.02 15.84 16.57
BELEBELE 66.52 21.22 38.35 56.61 64.33 59.82 66.84
SIB200 40.89 11.23 21.65 33.37 42.20 34.07 44.94
OPUS tg→en 19.59 10.67 20.01 22.68 21.63 24.52 24.34
OPUS en→tg 13.32 7.71 13.04 16.33 13.70 16.93 17.23
Avg. 29.56 11.88 21.81 28.41 30.05 29.89 33.02
General Performance
ARC challenge 70.90 20.40 32.89 59.62 70.57 63.74 70.62
ARC easy 85.79 23.86 47.02 77.00 84.78 80.81 84.74
Hellaswag 40.38 13.48 22.53 34.99 39.80 34.90 40.21
PIQA 71.87 34.85 48.21 64.64 74.15 69.70 72.61
Winogrande 51.47 38.06 48.45 50.88 52.07 51.27 52.35
MMLU 58.07 22.32 32.42 49.47 57.83 52.49 57.52
Avg. 63.08 25.50 38.59 56.10 63.20 58.82 63.01
Table 2: Average results across 12 languages for Llama-3.1-8B. green: best; grey: second-best. Per-language and -task results for all models are in Appendix D.
5.1 Subnetwork Identification
Table 1 summarizes the number of language-specific neurons, corresponding trainable parameters, and the fraction of the model updated for each subnetwork within Llama-3.1-8B. Subnetwork sizes vary considerably: Welsh and Nepali subnetworks cover over 0.7% of model parameters, whereas Georgian and Macedonian subnetworks account for less than 0.3%. This variation might reflect differences in language complexity and the amount of language-specific signal captured by GlotCC. Subnetworks for Mistral-Nemo-12B and Aya-Expanse-8B (Appendix B) exhibit similar distributions and relative sizes.
Figure 2 and Appendix B.1 depict the distribution of the subnetwork neurons across layers for all 12 languages. We observe that most language-specific neurons are concentrated in the upper layers, and related languages exhibit moderate overlap in their neuron sets (such as Slovak and Slovenian, or Latvian and Lithuanian–see Appendix B.2 for details). These patterns align with prior findings for more high-resource languages from Gurgurov et al. (2025a). Overall, the weights associated with these subnetworks account for roughly 0.2% to 1% of the total model parameters, highlighting the efficiency of targeted adaptation.
5.2 Subnetwork Fine-tuning
5.2.1 Average Task Performance
Table 2 reports average downstream task performance across the 12 target languages for different fine-tuning configurations, along with results on general-purpose benchmarks for Llama-3.1-8B. Targeted sparse fine-tuning of language-specific subnetworks consistently yields the largest gains on target-language tasks, outperforming full-model, FFN-only, LoRA, IA3, and random-subset baselines, while fully preserving general capabilities as measured by MMLU and commonsense benchmarks (63.0 vs. 63.1 for the base model). For example, on FLORES, targeted subnetworks achieve average improvements of up to 5 BLEU points. This contrasts strongly with full-model and FFN-only fine-tuning, which not only fail to improve target-language performance but also substantially degrade general capabilities, a clear indicator of catastrophic forgetting and overfitting to the training distribution. The random neuron-subset and LoRA baselines show intermediate behavior: modest target-language gains but noticeable reductions in general performance, underscoring the benefit of principled subnetwork selection over arbitrary sparsity or auxiliary modules. IA3 does not degrade overall performance, but its improvements on target-language tasks are consistently small compared to targeted subnetwork fine-tuning.
5.2.2 Language-Wise Variation
As shown in Figure 3, the magnitude of improvement varies across languages. Languages such as Maltese and Estonian benefit the most from targeted fine-tuning, with average gains exceeding 8 points across all tasks, whereas Lithuanian and Slovak see only marginal improvements. A likely explanation is that higher-resource languages are already well represented in pre-training, limiting the marginal utility of additional fine-tuning data. In some cases, the model may even have encountered overlapping material during pre-training, reducing the benefit of adaptation (Lee et al., 2022b). By contrast, lower-resource languages, where the base model may have seen little to no data, gain substantially from subnetwork adaptation. This pattern is reflected in the moderate negative correlation between improvement scores and language resource availability (Appendix G), with improvements decreasing as resource levels increase.
The three models also exhibit distinct patterns in response to fine-tuning. Aya-Expanse-8B shows the most consistent improvements, with gains ranging from +2.1 (Slovak) to +21.5 (Maltese) and no negative results. Llama-3.1-8B achieves moderate gains across most languages (+1.7 to +8.8 points) but shows a small decrease for Georgian (-1.1). Mistral-Nemo-12B, despite its larger size, exhibits more variable behavior: strong gains for Maltese (+17.7) and Welsh (+10.4) are offset by declines for Latvian (-0.2), Nepali (-1.8), and Slovenian (-2.4). This pattern likely reflects differences in pre-training data composition and language capacity allocation. Notably, Aya-Expanse-8B–designed for multilingual capabilities–benefits most uniformly, suggesting models with stronger multilingual foundations are more amenable to subnetwork adaptation for low-resource languages.
The only observed decreases are for Georgian with Llama and Slovenian and Nepali with Mistral. Since both models still improve on these languages overall, and individual task results are mostly positive, these decreases may reflect that the allocated subnetwork is relatively small for effectively capturing language-specific knowledge for these languages, even if the absolute number of parameters is large. This suggests that some languages might benefit from different $\rho$% thresholds or alternative allocations across FFN components. Component-wise ablations (Appendix F) show that while fine-tuning all FFN projections performs best on average, some languages improve more by updating only specific components. We leave a systematic investigation of optimal subnetwork sizes and component-wise effects across languages to future work.
(a) Llama-3.1-8B
(b) Mistral-Nemo-12B
(c) Aya-Expanse-8B
Figure 3: Average improvement scores across all tasks for the target fine-tuning for each of the 12 languages.
5.2.3 Training Dynamics
Figure 13(a) illustrates the training dynamics for Afrikaans. Sparse fine-tuning of target-language subnetworks generally converges faster and reaches lower validation loss than random-subset, FFN-only, full-model, LoRA, or IA3 fine-tuning. While full-model and FFN-only fine-tuning occasionally achieve lower loss on the target language (see Appendix C) due to the greater capacity of the updated networks, this is often accompanied by pronounced overfitting and a substantial drop in general capabilities.
5.2.4 Effect of Training Data Size
Figure 4 further examines the effect of varying training data size (50M, 100M, 150M, and 200M tokens) for a subset of seven languages (Macedonian, Latvian, Nepali, Slovak, Afrikaans, Maltese, and Lithuanian) using Llama-3.1-8B. For five out of seven languages, we observe that increasing data leads to improved performance, suggesting that additional data could further enhance results. This differs from the findings of Zhao et al. (2024), who fine-tune an ablation-identified subset of neurons using between 100 and 800 documents and report saturation after 400 documents. However, for Lithuanian and Maltese we similarly observe a decrease with larger training token sizes, which may indicate that the additional data is not helpful (Lee et al., 2022b) or the capacity of the identified subnetwork is insufficient to effectively utilize it. Due to computational constraints and the limited availability of data for some languages (e.g., Maltese and Afrikaans, with at most 100–150M tokens), we leave further exploration to future work.
Figure 4: Performance scores against various data sizes for a set of languages when sparsely fine-tuning Llama-3.1-8B.
6 Effects of Sparse Fine-Tuning
Beyond performance gains, we investigate the mechanisms underlying successful language adaptation. We analyze two aspects: weight changes induced by sparse fine-tuning and the evolution of cross-lingual representations.
(a) Up Projection
(b) Gate Projection
(c) Down Projection
Figure 5: Per-layer changes in weight values for three considered weight types for Nepali in Llama-3.1-8B. The plots for the other 11 languages follow the same patterns and will be made available in our GitHub repository.
(a) Base
(b) Sparse Fine-tuning
(c) Average cosine similarity
Figure 6: Cosine similarity for residual-stream activations of the original and sparsely fine-tuned Llama-3.1-8B between Maltese and reference languages (left, middle) and average across target languages (right). Full results in Appendix H.
6.1 Weight Change Analysis
Our analysis reveals a notable asymmetry between weight types within FFN layers. As shown in Figure 5, absolute weight changes for gate_proj and up_proj remain small and tightly clustered around zero across all layers, indicating highly precise adjustments. In contrast, down_proj exhibits substantially larger weight changes, particularly in later layers, where language-specific neurons are most concentrated.
[…]
6.2 Representation Drift Analysis
To assess how targeted neuron fine-tuning affects cross-lingual representations, we measure whether parallel sentences become closer in embedding space after adaptation. Specifically, we compute cosine similarities between post-residual hidden states (mean-pooled across tokens) for Flores-200 sentences in the 12 target languages and 9 typologically diverse high-resource languages (Arabic, Chinese, English, French, German, Italian, Japanese, Russian, Spanish) across all layers (Figure 6). Results show consistent increases in cosine similarity after adaptation, particularly in later layers where early-layer representations already exhibit high similarity. As illustrated in Figure 6(c), targeted sparse fine-tuning achieves the strongest cross-lingual alignment (average cosine similarity of 90.5), outperforming all baselines including LoRA (87.5) and random-subset fine-tuning (87.7). Full fine-tuning, despite updating all parameters, achieves weaker alignment (87.3), likely due to representational drift from overfitting.
These findings indicate that selectively updating identified neurons produces more language-agnostic representations in later layers, improving cross-lingual alignment without compromising model structure. This aligns with concurrent work showing improved alignment when manually manipulating language-relevant activations (Sundar et al., 2025), and suggests that sparse, targeted adaptation may be preferable to dense fine-tuning even when cross-lingual transfer is the primary goal.
7 Conclusion and Future Work
We present a framework for enhancing LLM capabilities in underrepresented languages via fine-tuning of language-specific subnetworks. Identifying language neurons and tuning their weights, our approach achieves substantial gains across 12 low- and mid-resource languages while preserving general-purpose performance. These results demonstrate that sparse, targeted adaptation can inject new language capabilities without catastrophic forgetting, addressing a key limitation of existing approaches. Mechanistic analyses reveal that adaptation concentrates in down-projection weights of later layers and improves cross-lingual alignment more effectively than dense fine-tuning. Our release of language neuron identifications for over 100 languages provides a practical resource for extending models to underrepresented languages. More broadly, our framework exemplifies actionable interpretability (Mosbach et al., 2024): we show that mechanistic insights can directly guide model enhancement, improving capabilities while preserving general performance. This highlights the importance of interpretability methods that enable optimization, not just post hoc analysis.
Looking forward, key directions include adaptively determining subnetwork size and composition based on language typology, data availability, and model scale, instead of fixed thresholds. Extending subnetwork adaptation to extremely low-resource languages and to other modalities such as vision and speech are also highly promising.
Impact Statement
This work advances the field of Machine Learning by enabling large language models to better handle underrepresented languages. Potential societal impacts include broader access to AI capabilities for speakers of low- and mid-resource languages and more equitable language technology, though risks such as misuse or biased outputs remain.
Acknowledgments
This research was supported by the German Federal Ministry of Research, Technology and Space (BMFTR) as part of the project TRAILS (01IW24005) and lorAI - Low Resource Artificial Intelligence, a project funded by the European Union under GA No.101136646.
Appendix
Appendix A Limitations
While our results show that sparse, language-specific subnetwork fine-tuning is effective across multiple models and languages, several limitations remain.
First, the effectiveness of targeted subnetwork adaptation depends on accurately identifying language-specific neurons. This process can be sensitive to the quality, coverage, and representativeness of the monolingual data used for neuron identification, particularly for languages with limited or noisy corpora.
Second, our experiments primarily focus on mid- and low-resource languages with sufficient post-training data. How well the method scales to extremely low-resource settings or to languages that are highly distant from those seen during pre-training remains an open question.
Third, we adopt a fixed selection threshold and do not systematically vary the capacity of language-specific subnetworks (i.e., the parameter $\rho$). Our results suggest that different languages may require different amounts of capacity to effectively capture language-specific knowledge, and that the interaction between subnetwork size and available training data warrants further study, potentially at the per-language or per-language-family level.
Finally, due to computational constraints, we did not run Full fine-tuning, FFN-only, LoRA, or IA3 baselines for Mistral-Nemo-12B and Aya-Expanse-8B. Based on our observations with Llama-3.1-8B, we expect these dense and PEFT methods to exhibit similar behavior on these models, but formal verification remains future work.
Appendix B Language Neuron Details
B.1 Language Subnetworks
Figures B.1 and 10 illustrate how language-specific neurons are distributed across layers for each model. While the absolute number of selected neurons varies substantially across languages and architectures, consistent structural patterns emerge. In particular, subnetworks tend to be sparse in early layers and increasingly concentrated in later layers, suggesting that language-specific specialization primarily occurs at higher levels of abstraction. Differences in subnetwork size across languages likely reflect a combination of pre-training exposure, typological similarity to high-resource languages, and model-specific capacity allocation.
Figure 7: Neuron distributions over the layers of Mistral-Nemo-12B for all 12 languages. Asterisks indicate languages with non-Latin scripts.
Lang. # Neurons # Trainable %
Afrikaans 3557 54.6M 0.45
Welsh 8239 126.6M 1.03
Estonian 3815 58.6M 0.48
Icelandic 5644 86.7M 0.71
Georgian 2028 31.1M 0.25
Lithuanian 3840 59.0M 0.48
Latvian 3741 57.5M 0.47
Macedonian 2403 36.9M 0.30
Maltese 3611 55.5M 0.45
Nepali 5567 85.5M 0.70
Slovak 2310 35.5M 0.29
Slovenian 2365 36.3M 0.30
Figure 8: Neuron and parameter counts for different language subnetworks, and percentages of the model being trained within Mistral-Nemo-12B.
Figure 9: Neuron distributions over the layers of Aya-Expanse-8B for all 12 languages. Asterisks indicate languages with non-Latin scripts.
Lang. # Neurons # Trainable %
Afrikaans 2091 25.7M 0.32
Welsh 5820 71.5M 0.88
Estonian 3628 44.6M 0.55
Icelandic 4303 52.9M 0.65
Georgian 2432 29.9M 0.37
Lithuanian 3443 42.3M 0.52
Latvian 3536 43.5M 0.54
Macedonian 2439 30.0M 0.37
Maltese 3201 39.3M 0.49
Nepali 3870 47.6M 0.59
Slovak 1924 23.6M 0.29
Slovenian 2538 31.2M 0.39
Figure 10: Neuron and parameter counts for different language subnetworks, and percentages of the model being trained within Aya-Expanse-8B.
B.2 Language Neuron Overlap
Figure 11 shows pairwise overlap between language-specific subnetworks, grouped by language family. Across models, related languages consistently exhibit higher overlap than unrelated ones, indicating partial sharing of language-relevant components rather than strictly disjoint representations.
(a) Llama-3.1-8B
(b) Mistral-Nemo-12B
(c) Aya-Expanse-8B
Figure 11: Language neuron overlap for all 12 languages across models.
B.3 Language Neuron Distributions
Below, we provide per-language distributions of identified language-specific neurons across layers for each model.
(a) Llama-3.1-8B
(b) Mistral-Nemo-12B
(c) Aya-Expanse-8B
Figure 12: Language neuron distributions for all 12 languages across models.
Appendix C Training Dynamics
This section presents validation loss curves for all 12 target languages, illustrating training dynamics across fine-tuning strategies.
(a) Afrikaans
(b) Welsh
(c) Estonian
(d) Icelandic
(e) Georgian
(f) Lithuanian
(g) Latvian
(h) Macedonian
(i) Maltese
(j) Nepali
(k) Slovak
(l) Slovenian
Figure 13: Validation loss curves for Llama-3.1-8B across languages.
Appendix D Downstream Results
In this appendix, we provide detailed per-language and per-task performance results for all models, complementing the averaged results reported in the main text.
D.1 Llama-3.1-8B
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE af 72.11 31.22 47.67 65.00 73.33 68.89 73.11
SIB200 af 35.29 27.45 30.39 31.37 42.16 45.59 44.12
[…]
Avg. 38.01 18.47 32.73 38.40 40.35 42.03 42.99
General Performance
arc challenge 70.90 25.75 35.45 64.55 73.91 62.21 71.24
arc easy 85.79 26.14 53.33 79.12 85.96 81.75 85.26
hellaswag 40.38 14.68 26.13 35.22 38.26 36.04 38.72
piqa 71.87 39.17 50.00 73.07 74.05 72.03 71.71
winogrande 51.38 49.17 50.43 50.91 52.49 47.04 51.46
mmlu 58.07 27.63 37.30 51.60 57.61 54.74 56.43
Avg. 63.07 30.42 42.11 59.08 63.71 58.97 62.47
Table 3: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
SIB200 cy 37.25 1.47 25.49 49.51 38.24 20.10 35.78
[…]
Avg. 25.43 17.95 27.00 31.33 25.95 26.61 28.65
General Performance
arc challenge 70.90 16.05 22.74 56.52 69.57 54.52 68.23
arc easy 85.79 16.84 30.70 74.56 85.61 76.14 84.04
hellaswag 40.38 12.31 25.27 29.93 38.66 30.88 40.79
piqa 71.87 3.97 49.46 58.98 73.72 66.97 68.39
winogrande 51.62 20.28 49.57 49.17 50.20 51.46 50.51
mmlu 58.07 22.27 25.67 45.79 58.26 48.14 56.69
Avg. 63.10 15.29 33.90 52.49 62.67 54.68 61.44
Table 4: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE et 66.67 15.00 42.00 56.22 64.00 58.11 65.00
SIB200 et 40.20 12.25 30.88 36.76 41.67 37.75 55.88
[…]
Avg. 26.57 8.60 21.55 27.81 27.17 28.99 33.30
General Performance
arc challenge 70.90 22.07 36.45 58.86 71.57 63.55 72.58
arc easy 85.79 21.40 55.09 77.89 84.91 81.58 84.74
hellaswag 40.38 13.92 22.14 36.11 39.84 37.06 41.82
piqa 71.87 40.10 54.35 59.63 75.14 72.96 75.24
winogrande 51.38 41.83 50.20 51.14 52.80 52.01 53.83
mmlu 58.07 20.64 33.90 48.66 58.59 52.84 57.35
Avg. 63.07 26.66 42.02 55.38 63.81 60.00 64.26
Table 5: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE is 62.67 21.44 41.67 50.89 58.00 51.00 66.11
SIB200 is 41.18 2.45 24.51 25.98 47.55 28.43 42.16
[…]
Avg. 26.53 11.25 21.06 23.89 27.03 24.22 29.41
General Performance
arc challenge 70.90 21.74 37.79 56.19 68.23 59.20 74.25
arc easy 85.79 23.33 55.96 74.39 83.68 76.84 85.61
hellaswag 40.38 8.29 25.24 32.65 40.81 27.75 44.67
piqa 71.87 26.55 54.79 69.48 76.01 61.48 77.31
winogrande 51.38 40.02 50.28 52.01 51.93 52.17 54.54
mmlu 58.07 17.18 35.86 46.77 57.09 45.40 58.92
Avg. 63.07 22.85 43.32 55.25 62.96 53.81 65.88
Table 6: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE ka 59.67 0.00 12.67 60.33 57.22 49.56 58.89
SIB200 ka 49.02 0.98 0.00 48.53 48.53 29.90 43.14
[…]
Avg. 24.16 1.88 5.64 24.04 23.75 19.27 23.05
General Performance
arc challenge 70.90 9.03 10.70 71.57 70.90 63.55 69.90
arc easy 85.79 10.18 15.44 85.26 85.96 82.98 84.56
hellaswag 40.38 1.24 6.33 42.39 42.30 33.13 41.56
piqa 71.87 12.13 8.81 73.67 72.69 64.53 71.16
winogrande 51.38 3.95 37.57 53.12 52.64 50.91 51.38
mmlu 58.07 4.51 18.42 58.00 58.00 55.26 56.69
Avg. 63.07 6.84 16.21 64.00 63.75 58.39 62.54
Table 7: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE lt 66.44 34.33 46.33 53.00 65.22 64.56 67.00
SIB200 lt 45.59 2.45 27.45 31.37 47.55 27.94 46.08
[…]
Avg. 29.63 11.54 21.95 26.10 29.93 28.12 31.50
General Performance
arc challenge 70.90 20.74 37.79 54.52 69.90 67.89 70.23
arc easy 85.79 27.89 59.12 74.04 84.56 80.53 85.26
hellaswag 40.38 20.49 29.02 31.45 41.50 38.14 41.37
piqa 71.87 39.28 51.63 57.56 74.37 69.15 74.21
winogrande 51.38 51.07 49.96 50.43 51.62 50.36 53.35
mmlu 58.07 27.24 36.64 45.85 58.20 52.19 57.02
Avg. 63.07 31.12 44.03 52.31 63.36 59.71 63.58
Table 8: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE lv 65.56 22.00 39.78 47.22 64.56 55.11 67.67
SIB200 lv 42.16 20.10 2.45 16.67 44.12 10.78 41.67
[…]
Avg. 29.62 13.01 19.13 24.10 30.24 25.02 32.20
General Performance
arc challenge 70.90 20.07 25.42 53.51 71.24 68.23 69.23
arc easy 85.79 27.54 36.67 70.70 84.21 81.58 84.56
hellaswag 40.38 14.48 17.34 31.23 40.71 35.03 41.73
piqa 71.87 39.50 43.69 55.93 75.46 71.82 74.59
winogrande 51.62 47.43 46.80 50.04 52.57 50.43 55.17
mmlu 58.07 24.89 30.70 45.59 58.07 53.43 59.63
Avg. 63.10 28.98 33.44 51.17 63.71 60.09 64.15
Table 9: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE mk 76.56 28.89 44.11 72.44 74.89 68.33 72.11
SIB200 mk 44.12 9.31 34.80 39.71 50.00 53.43 40.20
[…]
Avg. 33.25 12.60 24.83 33.98 34.56 36.89 34.98
General Performance
arc challenge 70.90 31.10 43.48 69.90 72.24 66.22 69.23
arc easy 85.79 39.82 64.39 84.39 85.61 82.11 85.61
hellaswag 40.38 26.41 25.06 37.91 41.80 32.57 36.43
piqa 71.87 49.67 53.26 73.72 75.73 68.44 67.36
winogrande 51.62 49.41 49.25 52.96 52.25 51.78 52.09
mmlu 58.07 30.96 39.65 56.69 58.39 54.67 57.48
Avg. 63.10 37.90 45.85 62.60 64.34 59.30 61.37
Table 10: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE mt 57.67 25.11 40.00 50.89 53.67 58.11 66.44
SIB200 mt 37.75 10.29 24.02 27.94 33.82 33.82 38.73
[…]
Avg. 30.22 14.37 30.13 27.34 30.42 34.93 39.01
General Performance
arc challenge 70.90 22.07 38.13 49.16 69.57 62.21 71.57
arc easy 85.79 21.23 56.84 72.46 83.51 82.11 83.86
hellaswag 40.38 9.70 26.72 31.45 36.30 35.07 39.00
piqa 71.87 42.11 50.87 61.97 70.73 72.85 73.34
winogrande 51.54 51.22 49.17 49.96 50.75 53.67 50.91
mmlu 58.07 26.26 32.66 45.00 55.91 52.58 57.87
Avg. 63.09 28.77 42.40 51.67 61.13 59.75 62.76
Table 11: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE ne 59.67 1.22 29.44 53.89 54.67 55.78 57.44
SIB200 ne 36.27 4.41 4.41 27.45 29.41 33.33 57.84
[…]
Avg. 26.89 4.74 12.93 26.54 26.14 27.45 32.65
General Performance
arc challenge 70.90 12.71 35.45 65.89 69.90 67.56 71.57
arc easy 85.79 19.47 49.12 81.75 82.46 82.98 84.39
hellaswag 40.38 16.38 14.32 37.46 36.73 36.68 36.31
piqa 71.87 36.83 51.25 62.35 74.59 73.83 71.11
winogrande 51.38 13.58 47.04 50.12 52.41 54.06 50.99
mmlu 58.07 17.50 33.18 54.08 57.67 54.08 57.74
Avg. 63.07 19.41 38.39 58.61 62.29 61.53 62.02
Table 12: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE sk 73.67 27.44 39.78 61.11 73.00 65.56 72.89
SIB200 sk 40.20 20.59 39.22 44.61 39.22 33.82 46.08
[…]
Avg. 33.48 16.86 26.49 32.76 33.61 32.44 35.56
General Performance
arc challenge 70.90 21.74 36.45 61.20 71.91 62.21 71.57
arc easy 85.79 23.86 51.05 78.25 86.49 78.60 85.44
hellaswag 40.38 15.04 27.87 40.04 40.76 36.74 43.68
piqa 71.87 43.25 54.30 59.25 72.91 70.29 71.76
winogrande 51.54 38.91 50.43 51.07 53.04 50.91 51.70
mmlu 58.07 21.36 35.99 52.25 58.20 51.53 57.67
Avg. 63.09 27.36 42.68 57.01 63.88 58.38 63.64
Table 13: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Llama-3.1-8B.
Metric Base Full FT FFN-Only FT LoRA IA3 Random Target
Monolingual Capabilities
[…]
BELEBELE sl 71.00 26.78 38.44 51.67 69.11 63.00 68.56
SIB200 sl 41.67 23.04 16.18 20.59 44.12 53.92 47.55
[…]
Avg. 30.88 11.32 18.31 24.66 31.41 32.75 32.90
General Performance
arc challenge 70.90 21.74 34.78 53.51 67.89 67.56 67.89
arc easy 85.79 28.60 36.49 71.23 84.39 82.46 83.51
hellaswag 40.38 8.79 24.92 34.08 39.98 39.72 36.48
piqa 71.87 45.65 56.15 70.13 74.43 72.03 75.08
winogrande 51.46 49.88 50.67 49.64 52.17 50.43 52.25
mmlu 58.07 27.43 29.13 43.31 57.94 55.06 56.69
Avg. 63.08 30.35 38.69 53.65 62.80 61.21 61.98
Table 14: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Llama-3.1-8B.
D.2 Mistral-Nemo-12B
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE af 79.67 75.56 79.44
SIB200 af 31.86 43.14 31.86
[…]
Avg. 35.11 39.93 38.57
General Performance
arc challenge 80.27 72.24 75.59
arc easy 89.82 83.33 87.72
hellaswag 40.86 38.31 42.32
piqa 80.03 77.15 77.20
winogrande 54.22 53.28 52.17
mmlu 61.66 55.85 61.33
Avg. 67.81 63.36 66.06
Table 15: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
SIB200 cy 35.78 31.37 40.20
[…]
Avg. 21.41 30.09 31.77
General Performance
arc challenge 80.27 63.88 71.57
arc easy 89.82 78.77 85.09
hellaswag 40.86 30.91 35.44
piqa 80.03 61.15 65.61
winogrande 54.22 51.07 51.93
mmlu 61.66 50.10 55.58
Avg. 67.81 55.98 60.87
Table 16: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE et 73.00 62.44 71.00
SIB200 et 32.35 29.90 36.76
[…]
Avg. 25.18 26.53 28.37
General Performance
arc challenge 80.27 71.91 74.92
arc easy 89.82 83.68 88.42
hellaswag 40.86 32.99 40.67
piqa 80.03 68.61 77.75
winogrande 54.22 52.09 53.35
mmlu 61.66 54.93 59.83
Avg. 67.81 60.70 65.82
Table 17: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE is 67.00 60.00 70.67
SIB200 is 34.80 25.98 38.24
[…]
Avg. 24.75 19.25 29.14
General Performance
arc challenge 80.27 64.55 76.59
arc easy 89.82 83.16 88.95
hellaswag 40.86 31.59 41.73
piqa 80.03 60.72 74.54
winogrande 54.22 51.70 53.67
mmlu 61.66 51.99 60.94
Avg. 67.81 57.28 66.07
Table 18: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE ka 71.89 68.56 72.33
SIB200 ka 30.39 37.25 32.35
[…]
Avg. 21.99 24.50 23.00
General Performance
arc challenge 80.27 69.23 77.26
arc easy 89.82 82.81 87.37
hellaswag 40.86 41.17 37.87
piqa 80.03 70.08 72.47
winogrande 54.22 52.33 53.99
mmlu 61.66 56.63 60.48
Avg. 67.81 62.04 64.91
Table 19: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lt 73.44 64.11 71.56
SIB200 lt 33.82 25.98 34.80
[…]
Avg. 26.55 25.46 28.61
General Performance
arc challenge 80.27 68.56 76.59
arc easy 89.82 86.32 88.77
hellaswag 40.85 36.47 41.65
piqa 80.03 67.25 72.20
winogrande 54.22 51.14 53.83
mmlu 61.66 52.25 60.16
Avg. 67.81 60.33 65.53
Table 20: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lv 72.89 60.44 66.00
SIB200 lv 34.80 46.08 30.88
[…]
Avg. 27.55 31.21 27.39
General Performance
arc challenge 80.27 64.21 72.91
arc easy 89.82 83.68 87.54
hellaswag 40.86 34.50 35.25
piqa 80.03 66.32 62.84
winogrande 54.22 51.22 51.93
mmlu 61.66 53.89 58.79
Avg. 67.81 58.97 61.54
Table 21: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mk 76.11 73.56 76.67
SIB200 mk 31.37 31.86 44.12
[…]
Avg. 29.82 32.59 32.47
General Performance
arc challenge 80.27 73.58 75.92
arc easy 89.82 87.02 87.89
hellaswag 40.86 37.81 44.30
piqa 80.03 74.97 76.17
winogrande 54.22 51.93 54.06
mmlu 61.66 57.28 59.57
Avg. 67.81 63.77 66.32
Table 22: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mt 41.33 64.67 72.44
SIB200 mt 30.88 44.61 33.33
[…]
Avg. 18.89 36.86 36.60
General Performance
arc challenge 80.27 68.23 74.92
arc easy 89.82 80.88 86.67
hellaswag 40.86 35.46 39.45
piqa 80.03 72.74 65.23
winogrande 54.22 51.54 52.57
mmlu 61.66 52.45 59.24
Avg. 67.81 60.22 63.01
Table 23: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE ne 64.11 51.56 60.67
SIB200 ne 34.31 12.25 22.55
[…]
Avg. 25.56 21.35 23.72
General Performance
arc challenge 80.27 51.51 72.58
arc easy 89.82 66.84 87.19
hellaswag 40.86 29.48 35.37
piqa 80.03 63.38 76.01
winogrande 54.22 52.49 53.12
mmlu 61.66 46.18 55.45
Avg. 67.81 51.65 63.29
Table 24: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE sk 79.78 71.89 77.33
SIB200 sk 34.80 44.12 28.43
[…]
Avg. 30.04 34.08 32.01
General Performance
arc challenge 80.27 70.90 75.59
arc easy 89.82 85.09 88.77
hellaswag 40.86 36.49 40.12
piqa 80.03 70.95 78.62
winogrande 54.22 51.62 55.49
mmlu 61.66 53.36 60.42
Avg. 67.81 61.40 66.50
Table 25: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Mistral-Nemo-12B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE sl 77.56 68.33 71.89
SIB200 sl 32.84 32.35 26.96
[…]
Avg. 27.36 26.64 24.93
General Performance
arc challenge 80.27 68.56 72.91
arc easy 89.82 83.51 84.74
hellaswag 40.86 34.75 37.04
piqa 80.03 64.58 75.24
winogrande 54.22 53.04 54.14
mmlu 61.66 55.13 57.41
Avg. 67.81 59.93 63.58
Table 26: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Mistral-Nemo-12B.
D.3 Aya-Expanse-8B
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE af 76.89 81.00 81.56
SIB200 af 42.16 49.51 64.22
[…]
Avg. 36.84 42.70 44.25
General Performance
arc challenge 77.26 78.26 75.59
arc easy 90.00 88.95 90.53
hellaswag 56.15 43.21 40.11
piqa 77.48 78.35 79.98
winogrande 55.72 54.22 53.43
mmlu 88.11 83.08 85.76
Avg. 74.12 71.01 70.90
Table 27: Evaluation results comparing different fine-tuning configurations across various tasks for Afrikaans for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
SIB200 cy 32.35 67.65 58.33
[…]
Avg. 17.13 29.64 28.63
General Performance
arc challenge 77.26 72.24 76.25
arc easy 90.00 87.72 88.25
hellaswag 56.15 38.89 55.82
piqa 77.48 77.64 81.23
winogrande 55.72 56.12 56.04
mmlu 88.11 77.14 83.80
Avg. 74.12 68.29 73.56
Table 28: Evaluation results comparing different fine-tuning configurations across various tasks for Welsh for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE et 54.00 71.78 68.78
SIB200 et 39.71 50.49 66.67
[…]
Avg. 20.93 33.36 33.61
General Performance
arc challenge 77.26 73.91 73.24
arc easy 90.00 89.12 87.54
hellaswag 56.15 49.79 44.84
piqa 77.48 76.17 80.47
winogrande 55.72 55.09 57.30
mmlu 88.11 82.04 84.78
Avg. 74.12 71.02 71.36
Table 29: Evaluation results comparing different fine-tuning configurations across various tasks for Estonian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE is 53.22 70.00 72.78
SIB200 is 33.33 59.80 62.75
[…]
Avg. 21.07 33.50 34.06
General Performance
arc challenge 77.26 75.59 77.59
arc easy 90.00 88.77 89.82
hellaswag 56.15 38.54 45.24
piqa 77.48 76.55 79.60
winogrande 55.72 54.22 53.75
mmlu 88.11 85.56 87.79
Avg. 74.12 69.87 72.30
Table 30: Evaluation results comparing different fine-tuning configurations across various tasks for Icelandic for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE ka 53.56 58.44 60.44
SIB200 ka 48.04 68.63 64.71
[…]
Avg. 21.74 27.08 26.90
General Performance
arc challenge 77.26 73.58 77.93
arc easy 90.00 88.07 89.82
hellaswag 56.15 44.69 52.88
piqa 77.48 81.56 81.56
winogrande 55.72 52.96 54.85
mmlu 88.11 84.00 87.07
Avg. 74.12 70.81 74.02
Table 31: Evaluation results comparing different fine-tuning configurations across various tasks for Georgian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lt 70.22 74.67 76.22
SIB200 lt 45.59 43.63 46.08
[…]
Avg. 28.56 30.45 33.32
General Performance
arc challenge 77.26 73.91 74.92
arc easy 90.00 86.84 87.54
hellaswag 56.15 46.50 48.36
piqa 77.48 78.78 80.69
winogrande 55.72 53.35 55.88
mmlu 88.11 82.95 86.28
Avg. 74.12 70.39 72.28
Table 32: Evaluation results comparing different fine-tuning configurations across various tasks for Lithuanian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE lv 56.56 77.33 75.78
SIB200 lv 42.16 62.75 67.16
[…]
Avg. 23.21 38.17 38.03
General Performance
arc challenge 77.26 69.23 77.93
arc easy 90.00 86.14 89.47
hellaswag 56.15 42.04 52.23
piqa 77.48 77.26 81.88
winogrande 55.72 55.09 54.85
mmlu 88.11 81.71 87.20
Avg. 74.12 68.58 73.93
Table 33: Evaluation results comparing different fine-tuning configurations across various tasks for Latvian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mk 67.33 77.22 78.67
SIB200 mk 43.14 66.67 70.10
[…]
Avg. 28.24 41.03 42.24
General Performance
arc challenge 77.26 75.59 78.60
arc easy 90.00 87.89 90.00
hellaswag 56.15 48.53 43.78
piqa 77.48 78.18 82.05
winogrande 55.72 55.56 54.30
mmlu 88.11 84.39 88.18
Avg. 74.12 71.69 72.82
Table 34: Evaluation results comparing different fine-tuning configurations across various tasks for Macedonian for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE mt 43.56 74.78 74.56
SIB200 mt 45.59 59.80 68.63
[…]
Avg. 22.97 43.76 44.47
General Performance
arc challenge 77.26 76.92 74.58
arc easy 90.00 87.89 87.89
hellaswag 56.15 39.55 42.98
piqa 77.48 79.27 79.76
winogrande 55.72 53.28 55.09
mmlu 88.11 84.32 85.50
Avg. 74.12 70.21 70.97
Table 35: Evaluation results comparing different fine-tuning configurations across various tasks for Maltese for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE ne 45.89 63.11 66.00
SIB200 ne 47.55 65.20 69.12
[…]
Avg. 23.96 34.03 34.87
General Performance
arc challenge 77.26 73.24 79.26
arc easy 90.00 87.72 90.88
hellaswag 56.15 41.11 49.52
piqa 77.48 81.28 82.64
winogrande 55.72 53.99 54.85
mmlu 88.11 85.50 88.96
Avg. 74.12 70.47 74.35
Table 36: Evaluation results comparing different fine-tuning configurations across various tasks for Nepali for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE sk 80.56 79.33 81.00
SIB200 sk 52.94 42.16 54.90
[…]
Avg. 36.92 36.28 38.99
General Performance
arc challenge 77.26 77.59 74.92
arc easy 90.00 87.72 88.77
hellaswag 56.15 52.13 48.54
piqa 77.48 79.38 79.16
winogrande 55.72 53.99 54.70
mmlu 88.11 83.87 86.02
Avg. 74.12 72.45 72.02
Table 37: Evaluation results comparing different fine-tuning configurations across various tasks for Slovak for Aya-Expanse-8B.
Metric Base Random Target
Monolingual Capabilities
[…]
BELEBELE sl 67.22 78.56 80.22
SIB200 sl 40.69 49.51 54.90
[…]
Avg. 27.05 34.50 36.00
General Performance
arc challenge 77.26 69.23 78.60
arc easy 90.00 85.79 89.82
hellaswag 56.15 47.88 51.96
piqa 77.48 76.61 79.87
winogrande 55.72 54.93 57.46
mmlu 88.11 80.54 86.81
Avg. 74.12 69.16 74.09
Table 38: Evaluation results comparing different fine-tuning configurations across various tasks for Slovenian for Aya-Expanse-8B.
Appendix E Generation Prompts
E.1 Task Prompts
We use standardized prompt templates for each task category to ensure consistency across models and datasets. Below we show the representative formats. Placeholders such as {source_text}, {passage}, or {question} are replaced with the actual content at runtime.
Translation.
Translate this sentence into {target_name}: {source_text} Translation:
Classification.
Classify the following text into one of these topics:
A) Science/Technology
B) Travel
C) Politics
D) Sports
E) Health
F) Entertainment
G) Geography
Text: {text}
Answer:
Comprehension.
Read the following passage and pick the right answer.
Passage: {passage}
Question: {question}
{options}
Answer:
General Capabilities.
We adopt dataset-specific templates for MMLU and commonsense reasoning benchmarks:
  • MMLU:
    Answer the following multiple choice question:
    {question}
    {options}
    Answer:
  • HellaSwag:
    Complete the following scenario by choosing the most likely continuation:
    {ctx}
    {options}
    Answer:
  • PIQA:
    Choose the most appropriate solution for the following goal:
    Goal: {goal}
    A) {sol1}
    B) {sol2}
    Answer:
  • Winogrande:
    Complete the sentence by choosing the correct option:
    {sentence}
    A) {option1}
    B) {option2}
    Answer:
  • ARC (Easy/Challenge):
    Answer the following question:
    {question}
    {options}
    Answer:
E.2 Generation Settings
The generation is done using the following parameters:
SamplingParams(
    temperature=0,
    repetition_penalty=1.1,
    max_tokens=128,
    top_p=1.0,
    top_k=-1,
    stop_token_ids=[eos_token_id] if eos_token_id is not None
            else [],
    stop=[
[…]
        "Question:", "Q:",
        "Answer:", "A:",
        "Translation:", "The translation", "Translate:",
        "Summary:", "The summary", "Summarize:",
        "Passage:", "The passage", "Text:",
        "Options:", "The options", "The choices", "Choices:",
        "Context:", "The context", "The problem", "Problem:",
[…]
        "</s>", "<|endoftext|>",
        "Human:", "Assistant:"
    skip_special_tokens=True
Appendix F FFN Component Ablations
To understand which parts of the FFN contribute most to language-specific adaptation, we perform ablations by fine-tuning individual or combinations of the up-, gate-, and down-projection matrices of the identified language subnetworks. Table 39 reports the resulting performance across a subset of six languages.
FFN component mt af sk mk lv ne Avg.
Full subnetwork size 37.0M 31.7M 23.8M 21.4M 39.5M 56.8M
base 30.04 37.80 33.43 33.32 29.88 27.06 31.92
ffn (up+gate+down) 39.01 42.99 35.56 34.98 32.20 32.65 36.23
up 37.17 39.98 33.48 35.27 30.83 30.69 34.57
gate+up 38.14 42.18 34.81 34.13 31.28 27.23 34.63
down 34.99 40.60 35.31 36.57 30.36 28.66 34.42
up+down 37.13 44.11 35.00 35.28 33.71 29.72 35.83
Table 39: Average performance of fine-tuning different FFN components of Llama-3.1-8B across six languages. up, gate, and down denote the identified language-specific components within the FFN up-, gate-, and down-projection weight matrices, respectively. Fine-tuning a single matrix corresponds to one third of the full language subnetwork size, while fine-tuning two matrices corresponds to two thirds.
Appendix G Correlation with Language Resource Availability
To examine whether the observed improvements from targeted fine-tuning are systematically related to language resource availability, we analyze correlations using two complementary metrics.
CC-100 Corpus Size.
We use the CC-100 corpus size (in MB) as a proxy for the relative digital resource availability across languages (Conneau et al., 2020). While these values represent web-crawled text from a specific snapshot, they serve as reasonable approximations of the relative proportions of available internet data per language. The ratios between languages remain informative even if absolute quantities have changed over time.
Joshi Resource Class.
We additionally employ the taxonomy proposed by Joshi et al. (2020), which classifies languages into six categories (0–5) based on both labeled and unlabeled data availability. This discrete classification considers multiple factors beyond raw corpus size, including the availability of annotated datasets and representation in NLP research. For our target languages, resource classes range from 1 (Maltese) to 3 (Estonian, Lithuanian, Latvian, Slovak, Slovenian).
LLaMA-3.1-8B Aya-Expanse-8B Mistral-NeMo-12B
Figure 14: Correlation between monolingual improvement scores and language resource metrics. Top row: CC-100 corpus size in MB, used as a proxy for relative internet data availability (Conneau et al., 2020). Bottom row: Resource class from the Joshi et al. taxonomy (Joshi et al., 2020). Pearson correlation coefficients are shown; * indicates $p<0.05$. Performance gains from targeted fine-tuning are highest for lower-resource languages and diminish as resource availability increases, suggesting limited marginal benefit for languages already well represented during pre-training.
Appendix H Cross-lingual Alignment
We additionally provide detailedper-language cosine similarities between parallel sentences before and after fine-tuning for all models. These results complement Section 5 by illustrating how targeted subnetwork adaptation affects cross-lingual alignment across languages.
ar de en es fr it ja ru zh
Baseline 85.55 92.88 90.63 91.47 91.21 91.45 86.55 88.21 87.46
Full 86.73 91.90 94.65 91.14 91.48 89.98 86.10 87.36 88.83
MLP 84.22 90.39 94.05 89.85 90.39 89.62 85.06 87.48 88.19
Adapter 84.79 95.11 92.10 93.23 93.21 93.07 86.58 88.59 87.47
Random 86.76 95.06 92.94 92.84 92.95 92.66 88.15 90.18 89.31
Target 87.75 96.10 95.03 94.24 94.52 93.89 90.24 91.25 91.46
Table 40: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Afrikaans) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 83.22 87.82 85.92 87.72 87.58 88.11 83.09 84.76 83.24
Full 88.11 90.77 92.58 91.70 91.88 92.55 87.44 87.37 89.85
MLP 84.88 90.89 91.98 89.65 90.63 90.65 85.80 87.55 88.96
Adapter 88.13 93.40 90.40 93.30 93.29 94.08 86.11 89.41 86.25
Random 89.94 92.71 89.41 92.48 92.39 92.13 86.22 89.49 86.88
Target 94.35 95.49 93.20 94.90 95.57 95.62 91.70 92.71 91.57
Table 41: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Welsh) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 85.45 91.29 89.18 90.46 90.19 90.67 86.09 88.40 86.75
Full 83.70 91.91 93.61 91.33 91.97 93.14 85.91 87.06 89.27
MLP 80.59 90.15 92.08 87.14 88.69 89.03 82.08 86.85 85.91
Adapter 86.19 93.87 90.52 92.56 92.51 92.83 86.99 89.58 87.47
Random 84.61 93.27 89.25 91.13 91.01 91.38 84.89 89.54 85.50
Target 90.07 95.43 93.87 94.13 93.91 93.73 90.77 92.25 91.84
Table 42: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Estonian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 82.82 88.54 85.85 87.65 87.45 87.89 83.22 85.24 83.55
Full 89.21 90.85 92.80 90.94 90.99 88.86 88.24 88.52 90.31
MLP 80.94 90.20 90.23 88.49 88.76 89.58 78.67 84.65 82.93
Adapter 85.49 93.79 89.50 92.61 92.24 92.57 85.22 88.89 86.21
Random 87.74 91.14 89.89 91.65 90.48 90.17 85.29 91.03 87.21
Target 91.14 93.44 93.11 94.09 94.10 94.25 90.11 91.98 90.65
Table 43: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Icelandic) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 66.64 68.23 61.90 66.87 67.40 68.37 66.58 70.69 64.28
Full 64.74 64.17 57.36 64.37 63.03 64.32 63.36 66.32 61.19
MLP 67.84 69.05 61.96 67.05 67.74 68.95 67.34 69.48 65.94
Adapter 65.52 67.76 61.55 66.33 66.81 67.67 65.41 70.47 63.65
Random 67.02 69.26 62.24 67.80 68.14 69.37 66.86 71.79 65.19
Target 69.54 70.67 64.23 69.01 69.66 70.64 69.20 73.41 67.32
Table 44: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Georgian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 83.79 89.55 86.95 89.12 88.88 89.56 84.84 87.59 85.20
Full 86.30 92.68 92.10 91.05 91.04 91.40 88.21 87.89 89.71
MLP 82.28 91.34 91.10 89.48 89.88 90.35 85.49 86.40 87.62
Adapter 83.45 91.75 88.31 91.17 90.83 91.67 84.16 88.73 84.43
Random 83.26 91.59 88.84 91.37 90.73 91.36 84.83 88.83 85.68
Target 86.99 93.23 91.85 92.78 92.50 92.47 88.56 91.26 90.07
Table 45: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Lithuanian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 82.73 88.36 85.37 87.88 87.58 88.25 83.37 86.37 83.76
Full 86.70 89.06 90.55 90.20 89.57 88.98 87.45 84.64 89.21
MLP 80.97 88.03 88.93 86.05 86.14 86.56 82.48 84.67 86.19
Adapter 84.62 92.14 88.45 91.85 91.36 92.08 84.51 89.09 86.13
Random 82.07 91.14 87.65 90.65 90.18 90.39 83.74 87.63 84.94
Target 86.60 93.48 91.73 93.19 93.04 93.14 88.61 90.79 89.61
Table 46: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Latvian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 87.36 87.33 86.16 87.00 86.85 87.38 86.00 92.75 86.12
Full 85.37 89.55 90.10 86.95 89.35 87.79 84.21 92.15 87.69
MLP 84.71 88.09 90.33 87.73 87.95 88.06 86.25 90.71 88.38
Adapter 87.53 86.66 85.75 86.49 86.15 86.97 84.81 97.54 85.33
Random 94.88 89.60 87.52 89.51 89.00 90.22 88.70 98.18 88.11
Target 92.74 92.92 92.01 92.99 92.63 93.00 91.52 98.72 92.00
Table 47: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Macedonian) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 84.92 87.62 85.09 87.37 87.32 89.50 82.94 84.95 83.17
Full 86.03 90.98 91.14 89.55 89.71 90.60 85.43 88.60 85.49
MLP 86.29 90.72 90.80 90.55 89.50 90.07 86.65 88.56 87.35
Adapter 88.20 93.48 89.95 93.85 93.66 96.17 85.00 88.92 85.74
Random 89.54 91.84 90.38 91.49 91.17 93.10 86.12 89.70 86.50
Target 92.23 93.53 92.42 92.79 92.63 94.90 89.06 90.80 89.52
Table 48: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Maltese) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 85.43 84.39 82.60 83.91 83.85 84.81 85.77 86.33 84.49
Full 88.65 88.61 89.58 89.29 89.63 89.65 88.12 89.33 89.39
MLP 83.47 83.90 86.27 84.17 84.67 83.77 85.09 85.00 86.68
Adapter 89.31 86.91 84.48 86.07 85.95 86.98 88.93 89.12 86.84
Random 92.35 87.82 84.56 87.06 86.96 87.95 90.76 91.25 88.24
Target 94.93 92.81 89.43 91.40 91.43 92.10 94.87 94.62 92.49
Table 49: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Nepali) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 87.04 91.90 90.14 91.61 91.24 91.61 87.26 90.57 87.85
Full 87.16 92.87 93.77 91.80 91.93 91.83 88.86 88.68 89.80
MLP 84.25 91.34 93.15 90.85 90.58 90.91 84.58 88.15 87.52
Adapter 85.71 92.33 90.44 92.23 91.83 92.12 86.20 90.18 87.07
Random 87.06 92.57 91.00 92.11 91.82 91.69 87.97 91.88 88.57
Target 87.56 93.88 93.35 93.52 93.25 93.33 89.86 92.60 90.45
Table 50: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Slovak) and other languages for Llama-3.1-8B.
ar de en es fr it ja ru zh
Baseline 85.88 91.34 89.69 91.13 90.40 91.31 86.13 89.15 86.98
Full 86.41 89.81 93.60 90.56 89.94 90.11 86.23 89.84 88.16
MLP 84.26 89.02 92.18 89.38 89.64 89.61 84.05 87.33 87.56
Adapter 86.40 92.67 91.10 92.64 92.03 92.60 86.11 90.97 87.25
Random 84.48 92.45 90.52 91.82 91.10 92.14 84.45 91.15 85.49
Target 88.57 94.18 93.79 94.25 93.47 94.17 88.72 93.27 90.21
Table 51: Layer-averaged Cosine similarity (expressed as percentages) between the target language (Slovenian) and other languages for Llama-3.1-8B.